### v1

# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

Abstract.
The Fu–Kane–Mele $\mathbb{Z}_{2}$ index characterizes two-dimensional time-reversal symmetric topological phases of matter. We shed some light on some features of this index by investigating projection-valued maps endowed with a fermionic time-reversal symmetry.
Our main contributions are threefold. First, we establish a decomposition theorem, proving that any such projection-valued map admits a splitting into two projection-valued maps that are related to each other via time-reversal symmetry. Second, we provide a complete homotopy classification theorem for these maps, thereby clarifying their topological structure. Third, by means of the previous analysis, we connect the Fu–Kane–Mele index to the Chern number of one of the factors in the previously-mentioned decomposition, which in turn allows to exhibit how the $\mathbb{Z}_{2}$-valued topological obstruction to constructing a periodic and smooth Bloch frame for the projection-valued map, measured by the Fu–Kane–Mele index, can be concentrated in a single pseudo-periodic Kramers pair.
Introduction
Since the experimental discovery of the quantum Hall effect in the 1980s [von1980new] and its subsequent topological interpretation [PhysRevLett.51.2167], topological phases of matter have remained a central topic in condensed matter physics and mathematical physics. The theoretical prediction of the quantum spin Hall effect [Kane_2005], the establishment of the “periodic table” of topological insulators and superconductors by Kitaev [Kitaev_2009], and the first experimental confirmations [Ko_nig_2007, Ando2013] in the early 2000s have directed attention towards Topological Insulators (TIs), materials realizing symmetry-protected topological phases.
Kitaev’s classification, derived from $K$-theoretical arguments, divides the problem into ten symmetry classes, corresponding to the so-called “tenfold way”, first formulated independently by Schnyder, Ryu, Furusaki, and Ludwig [Schnyder_2008]. This framework provides a unifying picture of topological phases characterized by different combinations of fundamental symmetries: time-reversal, particle-hole, and chiral. Among these, the class AII—comprising systems with fermionic time-reversal symmetry—has attracted particular attention due to its rich topological structure and its physical realizations in two- and three-dimensional topological insulators.
Class AII corresponds to systems endowed with an antiunitary time-reversal symmetry $\Theta$ satisfying $\Theta^{2}=-1$. For crystalline and insulating systems, described by a periodic Hamiltonian with a spectral gap, one associates the so-called Fermi projector $P$, which defines the occupied subspace. This, in turn, gives rise to the Bloch bundle $E\to\mathbb{T}^{d}$ over the Brillouin torus $\mathbb{T}^{d}$, carrying an induced antiunitary equivariant structure. Fiber bundles subject to this type of symmetry are called “Quaternionic”-vector bundles, and have became object of independent geometric interest, beyond the specific case considered in this paper, due to their peculiar topological structure (see, e.g., [de2015classification, bunk2019topologicalinsulatorskanemeleinvariant]).
As shown in [panati2007triviality, monaco2015symmetry], the presence of $\Theta$ guarantees the existence of a global smooth trivialization of $E$. However, such a trivialization need not respect the time-reversal symmetry itself, and the obstruction to the existence of a symmetric trivialization carries non-trivial topological content.
In two-dimensional systems, this obstruction is quantified by the Fu–Kane–Mele (FKM) invariant, first introduced in [kane2005z]. While the original formulation of this invariant was already deeply topological in nature, several alternative formulations and interpretations have since been proposed, both in physical [FuKane2006, MooreBalents2007] and in mathematical terms, see e.g. [graf2013bulk, de2015classification, fiorenza2016z, monaco2015symmetry]. For an overview of the relation among some of the different approaches, we refer the reader to [Cornean_2017].
Following the approach initiated in [fiorenza2016z], the $\mathbb{Z}_{2}$ invariant can be interpreted as the obstruction to the existence of a globally symmetric trivialization of the Bloch bundle. In position space, the analog of a globally smooth trivialization of the Bloch bundle is the existence of exponentially localized Wannier bases. Such bases have a clear physical interpretation as representing localized electronic orbitals, and their existence is intimately related to the insulating nature of the system. The requirement that a Wannier basis be symmetric under time-reversal symmetry allows to implement the symmetry at the level of such orbitals, which can be useful when these are used, for example, to derive symmetric tight-binding models from continuum Hamiltonians [marzari2012maximally].
If, in addition to $\Theta$, the system possesses a unitary involutive symmetry $S$ commuting suitably with $\Theta$—corresponding, for example, to spin conservation—then the topological invariant persists and admits a simple physical interpretation in position space as the spin Chern number [Sheng_2006, Prodan_2009]. In this case, $S$ splits the Fermi projector into two components,
(1) $P_{\pm}:=\frac{1}{2}\big(P\pm PS\big),$
which are exchanged by $\Theta$, and the topological invariant corresponds to the Chern number of either component. As shown by Prodan [Prodan_2009], the spin Chern number is stable under small perturbations, remaining well-defined modulo $2$ even when the symmetry $S$ is slightly broken—thus capturing the $\mathbb{Z}_{2}$ topology of class AII systems.
In the present work, we prove that, for periodic systems endowed with fermionic time-reversal symmetry, it is always possible to decompose the Fermi projector $P$ into a pair of projectors $P_{+}$ and $P_{-}$, related to each other under time reversal symmetry, in direct analogy with the spin-conserving case. We show that the topological characterization of such systems is captured by the Chern number, modulo $2$, of one component of the decomposition. The $\mathbb{Z}_{2}$ nature of the invariant arises from the non-uniqueness of this decomposition. This result provides a new perspective on the geometric and physical meaning of the Fu–Kane–Mele index: unlike other topological formulations, it admits a clear position-space counterpart, namely the Chern marker of one of the decomposition projectors. This position-space counterpart of the invariant paves the way for extending several results previously obtained in the spin-conserving case to the more general symmetry class AII, such as the Středa formula [Monaco_2020].
The paper is structured as follows. The first section is devoted to describing the standard procedure by which, starting from a symmetric and periodic Hamiltonian, one obtains the object at the center of our mathematical study: periodic projection-valued maps. The second section revisits parallel transport and the first Chern number in the context of continuous, rather than differentiable, projection-valued maps. Although this section does not contain original results, we believe it is useful to provide a concise and coherent presentation of these techniques.
The third section introduces our definition of the $\mathbb{Z}_{2}$-valued topological invariant, which differs from previous definitions but, at first glance, offers no computational or interpretative advantage. The fourth section marks our first original and physically relevant contribution, as it establishes an equivalence between the previously defined invariant and the possibility of decomposing the considered projection-valued map into a pair of symmetric maps whose Chern numbers have parity consistent with the invariant.
In the fifth section, we study the problem of constructing sections for a projection-valued map. In particular, this problem was previously approached by interpreting the $\mathbb{Z}_{2}$ invariant as an obstruction to realizing a periodic, continuous, and symmetric section. Our approach, instead, interprets the invariant as a set of gluing conditions for a section that need not be periodic.
The sixth and final section confirms the completeness of the invariant with respect to unitary equivalence and homotopy relations. The article concludes with an appendix presenting a series of classical results in algebraic topology that are used throughout the paper.
1. Setting of the problem: from periodic Hamiltonians to projection-valued maps
In this section, we outline the standard procedure that allows one to associate a projection-valued map to a periodic Hamiltonian, following the construction reviewed in [panati2007triviality, panati2013bloch] to which we refer the reader for further details. We consider a quantum particle moving in a condensed matter system, whose configuration space $X\subset\mathbb{R}^{d}$ exhibits a periodic structure with respect to a lattice $\Lambda\simeq\mathbb{Z}^{d}$, commonly referred to as the Bravais lattice. Explicitly, this periodicity implies $X=\lambda+X$ for every $\lambda\in\Lambda$. Two types of models are typically used in this context: continuum models, in which $X=\mathbb{R}^{d}$, and tight-binding models, where $X$ is replaced by the discrete set of atomic positions and the quotient $X/\Lambda$ is finite. When additional internal degrees of freedom, such as spin, are present, the configuration space is naturally extended to $X\times\{1,\ldots,n\}$, and the corresponding Hamiltonian $H$ describing the system acts on $L^{2}(X)\otimes\mathbb{C}^{n}$. Furthermore, the system is assumed to be endowed with a fermionic time-reversal symmetry, i.e., an anti-unitary operator $\Theta$ acting on $L^{2}(X)\otimes\mathbb{C}^{n}$ such that $\Theta^{2}=-\mathds{1}$, commuting with the Hamiltonian $H$.
For the sake of presentation, in the following we focus on continuum models ($X=\mathbb{R}^{d}$), having in mind time-reversal symmetric systems modelled by Hamiltonians with spin-orbit coupling of Rashba type [rashba2005spin], but all that will be discussed applies just as well to tight-binding models, like the paradigmatic Kane–Mele Hamiltonian [Kane_2005, kane2005z].
The periodicity of the system is implemented at the operator level as a commutation relation between $H$ and lattice translations,
$HT_{\lambda}=T_{\lambda}H,$
where $(T_{\lambda}\phi)(x)=\phi(x-\lambda)$ for all $\lambda\in\Lambda$. Correspondingly, the fermionic time-reversal symmetry operator $\Theta$ is required to be itself translation-invariant, i.e., to commute with the lattice translations $T_{\lambda}$.
A powerful tool to study the spectral properties of such periodic Hamiltonians is the Bloch-Floquet-Zak transform [kuchment2016overview], which allows one to decompose the Hilbert space into fibers labeled by the quasi-momentum $k$. For any sufficiently rapidly decaying $\phi\in L^{2}(\mathbb{R}^{d})$, the transform is defined as the following function depending on $k\in\mathbb{R}^{d}$ and $y\in\mathbb{R}^{d}$:
$(\mathcal{U}_{\mathrm{BFZ}}\phi)_{k}(y)=\sum_{\lambda\in\Lambda}e^{-ik\cdot(\lambda-y)}(T_{\lambda}\phi)(y).$
Let us denote by $\Lambda^{*}=\{\lambda^{*}\in\mathbb{R}^{d}\mid\lambda^{*}\cdot\lambda\in 2\pi\mathbb{Z},\forall\lambda\in\Lambda\}$ the dual lattice, and by $\mathbb{B}$ a connected periodicity cell, the Brillouin zone, for $\Lambda^{*}$. With this choice, $(\mathcal{U}_{\mathrm{BFZ}}\phi)_{k}(y)$ naturally belongs to a fiber Hilbert space with constant fibers:
$(\mathcal{U}_{\mathrm{BFZ}}\phi)_{k}(y)\in\left\{f\in L^{2}_{\rm loc}(\mathbb{R}^{d},dk,\mathcal{H}_{f})\ \middle|\ f(k+\lambda^{*})=\tau_{\lambda^{*}}f(k)\right\}\simeq\int_{k\in\mathbb{B}}^{\oplus}dk\,\mathcal{H}_{f},$
where $\mathcal{H}_{f}$ is the space of $\Lambda$-periodic, locally square-integrable functions on $\mathbb{R}^{d}$, and $\tau_{\lambda^{*}}:\mathcal{H}_{f}\to\mathcal{H}_{f}$ acts as $\tau_{\lambda^{*}}(g)(y)=e^{i\lambda^{*}\cdot y}g(y)$. By also choosing a periodicity cell for $\Lambda$, the Wigner-Seitz cell $\mathbb{W}$, one can further identify $\mathcal{H}_{f}$ with $L^{2}(\mathbb{W})$. A classical result, which can be traced back to [bloch1929quantenmechanik, zak1964magnetic], reads:
Proposition 1.1.
The operator $\mathcal{U}_{\mathrm{BFZ}}$ extends uniquely to a unitary operator
$\mathcal{U}_{\mathrm{BFZ}}:L^{2}(\mathbb{R}^{d})\longrightarrow\int_{k\in\mathbb{B}}^{\oplus}dk\,L^{2}_{\rm per}(\mathbb{W}).$
When $H$ acts on $L^{2}(\mathbb{R}^{d})\otimes\mathbb{C}^{n}$, the transformed operator $(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})H(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})^{-1}$, acting on
$\mathcal{H}=\left(\int_{k\in\mathbb{B}}^{\oplus}dk\,L^{2}_{\rm per}(\mathbb{W})\right)\otimes\mathbb{C}^{n}\simeq\int_{k\in\mathbb{B}}^{\oplus}dk\,\left(L^{2}_{\rm per}(\mathbb{W})\otimes\mathbb{C}^{n}\right),$
is decomposable. More precisely, there exists a family of self-adjoint fiber operators $H(k)$ such that
$\big[(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})H(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})^{-1}f\big]_{k}=H(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{R}^{d},$
and these fibers satisfy the $\tau$-covariance condition
(2) $H(k+\lambda^{*})=\tau_{\lambda^{*}}H(k)\tau_{\lambda^{*}}^{-1},\quad\forall k\in\mathbb{R}^{d},\ \lambda^{*}\in\Lambda^{*}.$
The same procedure can be applied to $\Theta$, since it commutes with the translation operators. Consequently, there exists a family of anti-unitary operators $T(k)$ acting on $\mathcal{H}$ such that
$\left[(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})\Theta(\mathcal{U}_{\mathrm{BFZ}}\otimes\mathds{1})^{-1}(f)\right]_{k}=T(k)f_{-k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{R}^{d}.$
The appearance of the minus sign in front of $k$ is a direct consequence of the anti-linearity of $\Theta$, and reflects also on the $\tau$-covariance condition:
$T(k+\lambda^{*})=\tau_{\lambda^{*}}T(k)\tau_{\lambda^{*}},\quad\forall k\in\mathbb{B},\ \lambda^{*}\in\Lambda^{*}.$
In most physical models, $\Theta$ acts only on the second factor of the tensor product $L^{2}(X)\otimes\mathbb{C}^{n}$, which implies that the fibers $T(k)$ are independent of $k$. Henceforth, we will assume $T(k)\equiv T$. Although it is possible to generalize the analysis to a fully $k$-dependent $\Theta$, this extension lies beyond the scope of the present work. Straightforward computations then show that the commutation relation $TH=HT$ at the level of fibers reads
(3) $TH(k)=H(-k)T,\quad\forall k\in\mathbb{R}^{d}.$
Let us return to the fiber Hamiltonians. In most physically relevant models, as functions of the quasi-momentum $k$, these constitute an analytic family of operators in $L^{2}(\mathcal{W})\otimes\mathbb{C}^{n}$ in the sense of Kato [kato2013perturbation], and the spectrum of the full Hamiltonian satisfies $\sigma(H)=\bigcup_{k\in\mathbb{R}^{d}}\sigma(H(k))$ [reed1978iv]. Typically, $\sigma(H)$ consists of disjoint closed intervals, referred to as Bloch bands, which can be used, as shown by the following proposition, to construct projection-valued maps, which will be the main object of study in the following sections.
Proposition 1.2 (Riesz formula).
Let $H(k)$ be a family of self-adjoint operators on a Hilbert space $\mathcal{H}$, and let $\gamma$ be a smooth, closed complex curve encircling exactly one Bloch band $\Omega$. Then
$P_{\Omega}(k)=\frac{i}{2\pi}\int_{\gamma}\left(H(k)-z\mathds{1}\right)^{-1}dz$
defines a family of projection-valued maps satisfying $P_{\Omega}(k)=\chi_{\Omega}(H(k))$, where $\chi_{\Omega}$ is the characteristic function of $\Omega$. Moreover, $P_{\Omega}(k)$ inherits the same regularity in $k$ as $(H(k)-z)^{-1}$.
Proof.
Observe that if $f(w)=\frac{i}{2\pi}\int_{\gamma}\frac{1}{w-z}dz$, then $P_{\Omega}(k)=f(H(k))$. The integral computes the winding number of $\gamma$ around $w$, which is 1 if $w$ lies inside $\gamma$ and 0 otherwise. Hence, for the interval $\Omega$ enclosed by $\gamma$, we have $f=\chi_{\Omega}$, and thus
$P_{\Omega}(k)=f(H(k))=\chi_{\Omega}(H(k)).$
The regularity statement follows immediately, since derivatives can be brought inside the integral. ∎
An immediate corollary is that, for a sufficiently regular hamiltonian $H(k)$, the rank of $P_{\Omega}(k)$ is constant, since $\operatorname{tr}(P_{\Omega}(k))=\dim\operatorname{Ran}(P_{\Omega}(k))\equiv m\in\mathbb{N}$ defines a continuous map into a discrete set. Moreover, this family of projectors is again $\tau$-covariant, and in the time reversal symmetric setting it inherits a symmetry constraint from Equation (3):
$P_{\Omega}(k+\lambda^{*})=\tau_{\lambda^{*}}P_{\Omega}(k)\tau_{\lambda^{*}}^{-1},\quad TP_{\Omega}(k)=P_{\Omega}(-k)T,\quad\forall k\in\mathbb{B},\ \lambda^{*}\in\Lambda^{*}.$
Although the $\tau$-covariance of this family of projections is essential for defining the topological invariants we aim to study, the associated pseudoperiodicity can be cumbersome to handle. To address this, we employ the following construction, first introduced in [cornean2016construction]. We begin by considering a lattice basis $\lambda_{1}^{*},\lambda_{2}^{*},\dots,\lambda_{d}^{*}$ for $\Lambda^{*}$ and the corresponding unitary operators $\tau_{\lambda_{j}^{*}}(\phi)(y)=e^{i\lambda_{j}^{*}\cdot y}\phi(y)$ acting on $\mathcal{H}$. Using the spectral theorem for unitary operators, we find $d$ self-adjoint operators $L_{1},\dots,L_{d}$ such that $\tau_{\lambda_{j}^{*}}=e^{iL_{j}}$. Since they commute with each other, they can be chosen to satisfy $L_{a},L_{b}]=0$ for all $a,b=1,\dots,d$.
The $\tau$-covariance of the fibers of $\Theta$ implies
$T(k+\lambda_{j}^{*})=\tau_{\lambda_{j}^{*}}T(k)\tau_{\lambda_{j}^{*}}^{-1}.$
If the fibers are constant, this reduces to $Te^{iL_{j}}=e^{-iL_{j}}T$, and by functional calculus one can choose $L_{j}$ such that $TL_{j}=L_{j}T$, so that
$Te^{iL_{j}k_{j}}=e^{-iL_{j}k_{j}}T,\quad\forall k_{j}\in\mathbb{R}.$
When these properties hold simultaneously, we can define a unitary-valued map $U:\mathbb{R}^{d}\to\mathcal{U}(\mathcal{H})$ as
$U(k_{1}^{*},\dots,k_{d}^{*})=\prod_{j=1}^{d}e^{iL_{j}k_{j}^{*}}.$
By construction, it satisfies
$TU(k_{1}^{*},\dots,k_{d}^{*})=U(-k_{1}^{*},\dots,-k_{d}^{*})T.$
If $(k_{1}^{*},\dots,k_{d}^{*})=k^{*}$ are the coordinates of $k\in\mathbb{R}^{d}$ in the chosen lattice basis, we define
$\tilde{P}_{\Omega}(k)=U(k_{1}^{*},\dots,k_{d}^{*})P_{\Omega}(k)U(k_{1}^{*},\dots,k_{d}^{*})^{-1}.$
It is immediate to check that this projection-valued map is now $\Lambda^{*}$-periodic in $k$ while still satisfying the symmetry constraint
$\tilde{P}_{\Omega}(k)T=T\tilde{P}_{\Omega}(-k).$
The choice of $L_{1},\dots,L_{d}$ is not unique: if another unitary-valued map $U^{\prime}:\mathbb{R}^{d}\to\mathcal{U}(\mathcal{H})$ has the same properties, then the corresponding projections $\tilde{P}_{\Omega}$ and $\hat{P}_{\Omega}$ are unitarily equivalent via the periodic unitary-valued map $U^{\prime}(k^{*})U(k^{*})^{-1}$:
$\hat{P}_{\Omega}(k)=U^{\prime}(k^{*})U(k^{*})^{-1}\tilde{P}_{\Omega}(k)\left(U^{\prime}(k^{*})U(k^{*})^{-1}\right)^{-1},$
with the time-reversal constraint
$TU^{\prime}(k^{*})U(k^{*})^{-1}=U^{\prime}(-k^{*})U(-k^{*})^{-1}T,\quad\forall k^{*}\in\mathbb{R}^{d}.$
With this construction, we obtain a regular projection-valued map $P:\mathbb{R}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ that is $\Lambda^{*}$-periodic in $k$. We may reinterpret it as a projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ since $\mathbb{R}^{d}/\Lambda^{*}\simeq\mathbb{T}^{d}$, preserving the symmetry constraint
$TP(k)=P(-k)T,\quad\forall k\in\mathbb{T}^{d}.$
A family of projections $P$ satisfying the above assumptions will be the central object of study in this article.
Remark 1.3.
Before giving a formal definition, we briefly consider again the discrete setting, where the Bloch-Floquet transform and the procedure to eliminate the $\tau$-covariance can be summarized in a single step. As we will see, we again obtain a projection-valued map that is periodic and satisfies a symmetry constraint analogous to the one above.
Let $X\subset\mathbb{R}^{d}$ be a discrete configuration space with a lattice $\Lambda\simeq\mathbb{Z}^{d}$ such that $X=X+\lambda$ for all $\lambda\in\Lambda$, and suppose $X/\Lambda$ is finite. We can then select a finite set of points $\mathbb{W}=\{x_{1},\dots,x_{l}\}\subset X$ such that $X=\mathbb{W}+\Lambda$. Let $H$ be an Hamiltonian acting in $L^{2}(X)\otimes\mathbb{C}^{n}$, which we rewrite as
$L^{2}(X)\otimes\mathbb{C}^{n}\simeq L^{2}(\mathbb{W}\times\Lambda)\otimes\mathbb{C}^{n}\simeq L^{2}(\Lambda)\otimes\mathbb{C}^{l\cdot n}\simeq L^{2}(\mathbb{Z}^{d})\otimes\mathbb{C}^{l\cdot n}.$
We denote by $\mathcal{W}$ the corresponding isomorphism, which depends on the choice of $\mathbb{W}$. Applying the Fourier transform
$\mathcal{F}:L^{2}(\mathbb{Z}^{d})\to L^{2}(\mathbb{T}^{d},dk,\mathbb{C}),\quad(\mathcal{F}f)(k)=\sum_{n\in\mathbb{Z}^{d}}\frac{1}{(2\pi)^{d/2}}e^{ik\cdot n}f(n),$
to the first factor, the operator
[…]
is decomposable. That is, there exists a family of self-adjoint matrices $H(k)$ on $\mathcal{H}=\mathbb{C}^{l\cdot n}$ such that
$\big[(\mathcal{F}\otimes\mathds{1})\mathcal{W}H\mathcal{W}^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}(f)\big]_{k}=H(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{T}^{d}.$
Here, periodicity in $k$ holds directly, which is why we write $k\in\mathbb{T}^{d}$. The same steps as in the continuum case yield a projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ satisfying
$TP(k)=P(-k)T,\quad\forall k\in\mathbb{T}^{d},$
assuming constant fibers of $\Theta$.
It is important to note that the choice of discrete Wigner-Seitz cell $\mathbb{W}=\{x_{1},\dots,x_{n}\}$ is not unique, and the ordering of points could in principle affect the topological properties of the model. If a different cell $\mathbb{W}^{\prime}=\{x_{1}^{\prime},\dots,x_{n}^{\prime}\}$ is chosen, one defines another isomorphism $\mathcal{W}^{\prime}$ and obtains a new fiber decomposition
$\big[(\mathcal{F}\otimes\mathds{1})\mathcal{W}^{\prime}H(\mathcal{W}^{\prime})^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}(f)\big]_{k}=H^{\prime}(k)f_{k},\quad\forall f\in\mathcal{H},\ k\in\mathbb{T}^{d}.$
The two decompositions are related by a fibered unitary equivalence
$(\mathcal{F}\otimes\mathds{1})\mathcal{W}^{\prime}\mathcal{W}^{-1}(\mathcal{F}\otimes\mathds{1})^{-1}=\int_{k\in\mathbb{T}^{d}}^{\oplus}U(k),$
where $U(k)$ acts on $\mathbb{C}^{l\cdot n}$, so that
$U(k)H(k)U(k)^{-1}=H^{\prime}(k),$
and likewise for the associated projection-valued maps $P_{\Omega}(k)$ and $P_{\Omega}^{\prime}(k)$. To preserve the symmetric structure, one must choose $\mathbb{W}$ such that
$U(k)T=TU(-k),\quad\forall k\in\mathbb{T}^{d}.$
In both continuum and discrete settings, we thus obtain the central object of study in this paper.
Definition 1.4 (time-reversal symmetric projection-valued map).
Let $\mathcal{H}$ be a Hilbert space, $T$ an anti-unitary operator with $T^{2}=-\mathds{1}$, $m$ the rank of the projections, and $d$ the physical dimension of the model. A continuous projection-valued map $P:\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ is time-reversal symmetric if
(4) $TP(k)=P(-k)T,\quad\forall k\in\mathbb{T}^{d}.$
Two such maps $P_{0},P_{1}$ are unitarily equivalent if there exists a continuous unitary-valued map $U:\mathbb{T}^{d}\to\mathcal{U}(\mathcal{H})$ such that
(5) $TU(k)=U(-k)T,\quad P_{1}(k)=U(k)P_{0}(k)U(k)^{-1},\quad\forall k\in\mathbb{T}^{d}.$
Two maps $P_{0},P_{1}$ are homotopic if there exists a continuous projection-valued map $P:[0,1]\times\mathbb{T}^{d}\to\operatorname{Proj}_{m}(\mathcal{H})$ such that
(6) $P(0,k)=P_{0}(k),\quad P(1,k)=P_{1}(k),\quad TP(t,k)=P(t,-k)T,\quad\forall t\in[0,1],\ k\in\mathbb{T}^{d}.$
Remark 1.5 (Continuous vs. Analytic).
Since the family of projectors $\{P(k)\}_{k\in\mathbb{R}^{d}}$ appearing in physical models is usually analytic in $k$, one may ask whether real-analytic Bloch frames can be constructed. Following the argument of Lemma 2.3 in [cornean2016construction], one can show that whenever a continuous Bloch frame exists, it can be modified to a real-analytic one while preserving the time-reversal symmetry. Hence, in the following, we will freely interchange continuous, differentiable, and analytic structures.
Remark 1.6 (Role of Unitary Equivalences and Homotopies).
It is important to emphasize that unitary equivalence is associated with the choice of dimerization in tight-binding models, or with the selection of a unitary-valued map that removes the $\tau$-covariance in continuum models. In contrast, homotopies correspond to continuous deformations of physical parameters in the original Hamiltonian that do not close the spectral gaps. Under such deformations, the topological properties of $P$ are preserved. Since the physical interpretations of these two mathematical notions differ, it is natural to treat them separately. Moreover, when the two conditions coincide mathematically, they establish an equivalence between distinct physical aspects of the system.
2. Preliminary results
In this section, we introduce two constructions that will play a central role throughout the text. The first serves as a substitute for parallel transport in the continuous setting and will be crucial for defining the invariant. The second provides an equivalent formulation of the Chern number, which we will later use to reinterpret the invariant through a suitable decomposition of the projector .
2.1. Continuous parallel transport
In this subsection, we investigate the problem of constructing a unitary-valued map $U:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ associated with a given projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$, such that
$P(k)=U(k)\,P(0)\,U(k)^{-1}.$
A standard approach to this problem relies on parallel transport (see, e.g., [Cornean_2017, monaco2023az]), which requires at least $C^{1}$ regularity of the map $k\mapsto P(k)$. Here, we present a method based on the Kato–Nagy intertwiner, which applies already in the merely continuous setting and thus provides a genuine generalization of parallel transport to lower regularity.
Theorem 2.1 (Kato–Nagy unitary equivalence).
If $P,Q\in\operatorname{Proj}_{n}(\mathcal{H})$ are two projections with $\|P-Q\|<1$, then
(7) $U_{P\to Q}=\left[QP+(\mathds{1}-Q)(\mathds{1}-P)\right]\left[\mathds{1}-(P-Q)^{2}\right]^{-1/2}$
is a unitary operator such that $Q=U_{P\to Q}PU_{P\to Q}^{-1}$.
This formula will be used extensively; its proof can be found in [kato2013perturbation].
Proposition 2.2 (Kato–Nagy 1-d construction: continuous parallel transport).
If $P:\mathbb{T}^{1}\to\operatorname{Proj}_{m}(\mathcal{H})$ is a continuous projection-valued map, then there exists a continuous unitary-valued map $\tilde{U}:\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ such that
$P(k)=\tilde{U}(k)P(0)\tilde{U}(k)^{-1}.$
If $P$ is time-reversal symmetric, it is possible to choose $\tilde{U}$ so that $T\tilde{U}(k)=\tilde{U}(-k)T$.
Proof.
Divide $0,\pi]$ into a finite number of intervals $I_{j}=[t_{j},t_{j+1}]$ with $t_{0}=0$, $t_{J}=\pi$, such that $\|P(t_{j})-P(t)\|<1$ for all $t\in[t_{j},t_{j+1}]$. For $t\in I_{j}$, define $U_{t_{j}}^{t}=U_{P(t_{j})\to P(t)}$ according to Theorem 2.1. The unitary equivalence between $P(0)$ and $P(t)$ is then
$U(t)=U_{t_{j}}^{t}U_{t_{j-1}}^{t_{j}}\cdots U_{t_{0}}^{t_{1}}.$
If $P$ is time-reversal symmetric, extend $U$ to $\pi,0]$ via U(t)=TU(t)TU(t)=TU(-t)T; otherwise, replicate the same procedure for $t\in[-\pi,0]$. In both cases, $U:[-\pi,\pi]\to\mathcal{U}(\mathcal{H})$ satisfies $P(t)=U(t)P(0)U(t)^{-1}$ for all $t$.
This $U$ is generally not periodic in $t$, so consider $U(-\pi)^{-1}U(\pi)$, which commutes with $P(0)$. If $P$ is symmetric, it also satisfies $TU(-\pi)^{-1}U(\pi)=\left[U(-\pi)^{-1}U(\pi)\right]^{-1}T$. By the spectral theorem, there exists a self-adjoint operator $L$ commuting with $P(0)$ such that $U(-\pi)^{-1}U(\pi)=e^{iL}$. If $P$ is symmetric, $L$ can also be chosen to commute with $T$. Finally, define
$\tilde{U}(t)=U(t)e^{-itL/2\pi}.$
This construction ensures that $\tilde{U}$ is periodic, continuous, intertwines $P(0)$ with $P(t)$, and respects the symmetry constraint when $P$ does. ∎
Proposition 2.3 (Kato–Nagy 2-d construction).
If $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ is a continuous projection-valued map, then there exists a continuous unitary-valued map $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$, periodic in the second component but not in the first, such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}.$
If $P$ is time-reversal symmetric, $U$ can be chosen such that $TU(t,k_{2})=U(-t,-k_{2})T$ for all $(t,k_{2})\in[-\pi,\pi]\times\mathbb{T}^{1}$.
Proof.
Divide $0,\pi]$ into a finite number of intervals $t_{j},t_{j+1}]$ with $t_{0}=0$, $t_{J}=2\pi$, such that $\|P(t_{j},k_{2})-P(t,k_{2})\|<1/2$ for all $(t,k_{2})\in[t_{j},t_{j+1}]\times\mathbb{T}^{1}$. Define $U_{t_{j}}^{t}(k_{2})=U_{P(t_{j},k_{2})\to P(t,k_{2})}$ via Theorem 2.1.
To define the unitary equivalence between $P(0,0)$ and $P(t,k_{2})$, first apply Proposition 2.2 to the 1-d projection-valued map $P(0,k_{2})$ to construct a periodic (and possibly symmetric) unitary-valued map $\tilde{U}(k_{2})$ with $P(0,k_{2})=\tilde{U}(k_{2})P(0,0)\tilde{U}(k_{2})^{-1}$. Then, for $t\in[t_{j},t_{j+1}]$, set
$U(t,k_{2})=U_{t_{j}}^{t}(k_{2})U_{t_{j-1}}^{t_{j}}(k_{2})\cdots U_{t_{0}}^{t_{1}}(k_{2})\tilde{U}(k_{2}).$
For $t\in[-\pi,0]$, either impose $U(t,k_{2})=TU(-t,-k_{2})T$ if $P$ is symmetric, or replicate the previous construction otherwise. In both cases, the claim follows. ∎
2.2. Chern number through the matching matrices
In the following sections, it will be essential to adopt an alternative formulation of the Chern number: compare [Cornean_2017].
Definition 2.4 (Chern number).
Let $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ be a continuous, not necessarily symmetric, projection-valued map, and let $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ be a continuous unitary-valued map, periodic in the second argument, such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}.$
Then $U(-\pi,k_{2})^{-1}U(\pi,k_{2})$ commutes with $P(0,0)$, so after choosing an orthonormal basis $\{v_{j}\}_{j=1,\dots,n}$ spanning $\operatorname{Ran}(P(0,0))$, we define unitary matrices $\alpha(k_{2})\in M_{n,n}(\mathbb{C})$, called matching matrices, depending periodically and continuously on $k_{2}$, via
$\alpha_{a,b}(k_{2})=\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\rangle.$
The Chern number is then the integer obtained as the winding number of $\det(\alpha)$ :
$\operatorname{Ch}(P)=[\det(\alpha(\cdot))].$
If $\dim(\operatorname{Ran}(P(k)))=\infty$, we set $\operatorname{Ch}(P)=0$.
In the remainder of this section, we derive from the previous definition some fundamental properties of the Chern number. Although these results are well known, we believe that presenting their proofs in the notation introduced above may help the reader become familiar with the techniques employed in the following sections.
Proposition 2.5.
The above definition is independent of the choice of the unitary operator $U$. Moreover, if two projection-valued maps are unitarily equivalent or homotopic, their Chern numbers coincide.
Proof.
Well-posedness.  First, we show that the Chern number does not depend on the choice of the orthonormal basis $\{v_{j}\}$. If $\{v_{j}^{\prime}\}$ is another basis, it induces a different $\alpha^{\prime}(k_{2})$, which is unitarily equivalent to $\alpha(k_{2})$ via $A_{a,b}=\langle v_{a},v_{b}^{\prime}\rangle$, so
$\alpha(k_{2})=A^{-1}\alpha^{\prime}(k_{2})A\implies\det(\alpha(k_{2}))=\det(\alpha^{\prime}(k_{2})).$
Next, consider two unitary-valued maps $U,V:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ satisfying
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}=V(t,k_{2})P(0,0)V(t,k_{2})^{-1}.$
Then $V(t,k_{2})^{-1}U(t,k_{2})$ commutes with $P(0,0)$, and defining
$\beta_{a,b}(t,k_{2})=\langle v_{a},V(t,k_{2})^{-1}U(t,k_{2})v_{b}\rangle,$
continuity in $t$ and periodicity in $k_{2}$ imply that $\det\beta(t,\cdot)]$ is constant in $t$. By Corollary A.5 and $\det(AB)=\det(A)\det(B)$, we obtain
$\det(\alpha(\cdot))]=[\det(V(-\pi,\cdot)^{-1}V(\pi,\cdot)|_{\operatorname{Ran}(P(0,0))})],$
proving independence of the Chern number on $U$.
Unitary invariance.  Let $V(k)$ implement a unitary equivalence between two projection-valued maps $P_{0}(k)$ and $P_{1}(k)$:
$P_{1}(k)=V(k)P_{0}(k)V(k)^{-1}.$
Given a Kato-Nagy unitary $U$ such that
$P_{0}(t,k_{2})=U(t,k_{2})P_{0}(0,0)U(t,k_{2})^{-1},$
we can construct a corresponding unitary for $P_{1}$:
$P_{1}(t,k_{2})=V(t,k_{2})U(t,k_{2})V(0,0)^{-1}P_{1}(0,0)V(0,0)U(t,k_{2})^{-1}V(t,k_{2})^{-1}.$
Choosing a basis $\{v_{j}^{1}\}$ of $\operatorname{Ran}(P_{1}(0,0))$ and defining
$\alpha_{1}]_{a,b}(k_{2})=\langle v_{a}^{1},V(0,0)U(-\pi,k_{2})^{-1}V(-\pi,k_{2})^{-1}V(\pi,k_{2})U(\pi,k_{2})V(0,0)^{-1}v_{b}^{1}\rangle,$
periodicity of $V$ in $k_{2}$ and the relation $v_{j}^{0}:=V(0,0)^{-1}v_{j}^{1}$ lead to
$\alpha_{1}]_{a,b}(k_{2})=\langle v_{a}^{0},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}^{0}\rangle.$
Hence,
$\operatorname{Ch}(P_{1})=[\det(\alpha_{1})]=[\det(\alpha_{0})]=\operatorname{Ch}(P_{0}).$
Homotopy invariance.  Consider a homotopy $P_{t}(k)$ of projection-valued maps. By compactness of $0,1]\times\mathbb{T}^{2}$ and continuity, we can partition $0,1]=\bigcup_{j=0}^{J-1}[t_{j},t_{j+1}]$ such that
$\|P_{t_{j}}(k)-P_{t}(k)\|\leq 1/2,\quad\forall t\in[t_{j},t_{j+1}],k\in\mathbb{T}^{2}.$
Then, for each $j$, the Kato-Nagy unitary
$U_{j}(k)=\big[P_{t_{j+1}}(k)P_{t_{j}}(k)+(\mathds{1}-P_{t_{j+1}}(k))(\mathds{1}-P_{t_{j}}(k))\big]\big[\mathds{1}-(P_{t_{j}}(k)-P_{t_{j+1}}(k))^{2}\big]^{-1/2}$
implements a unitary equivalence between $P_{t_{j}}$ and $P_{t_{j+1}}$. By the previous item, Chern numbers are preserved along each step, yielding
$\operatorname{Ch}(P_{0})=\operatorname{Ch}(P_{t_{1}})=\dots=\operatorname{Ch}(P_{1}).\qed$
Proposition 2.6.
The Chern number is additive: if $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ and $Q:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$ satisfy $PQ\equiv 0$, then for $P+Q:\mathbb{T}^{2}\to\operatorname{Proj}_{n+m}(\mathcal{H})$,
$\operatorname{Ch}(P+Q)=\operatorname{Ch}(P)+\operatorname{Ch}(Q).$
Proof.
Applying Proposition 2.3 to $P$, $Q$, and $P+Q$, we obtain unitary equivalences
$U,V,Y:[0,2\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$
such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1},$
$Q(t,k_{2})=V(t,k_{2})Q(0,0)V(t,k_{2})^{-1},$
$(P+Q)(t,k_{2})=Y(t,k_{2})(P+Q)(0,0)Y(t,k_{2})^{-1}.$
Then we extend the operator $X:[0,2\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ by linearity with
$X(u)=\begin{cases}U(u)\quad\mbox{if}\quad Pu=u\\ V(u)\quad\mbox{if}\quad Qu=u\\ Y(u)\quad\mbox{if}\quad Pu=Qu=0\end{cases}$
to obtain a unitary-valued map with
$(P+Q)(t,k_{2})=X(t,k_{2})(P+Q)(0,0)X(t,k_{2})^{-1}.$
Now we compute the Chern number of $P+Q$ using a basis $\{v_{j}\}_{j\in\{1,\cdots,m+n\}}$ where the first $n$ span $P(0,0)$ and the other span $Q(0,0)$. In this way we have
$\alpha(k_{2})=\begin{pmatrix}\beta(k_{2})&0\\ 0&\gamma(k_{2})\end{pmatrix}$
where $\beta:\mathbb{T}^{2}\to U(n)$ and $\gamma:\mathbb{T}^{2}\to U(m)$ are such that
$\beta(k_{2})]_{a,b}=\left\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\right\rangle\quad\text{and}\quad[\gamma(k_{2})]_{a,b}=\left\langle v_{n+a},V(-\pi,k_{2})^{-1}V(\pi,k_{2})v_{n+b}\right\rangle\,.$
So using Corollary A.5 we obtain the thesis:
$\operatorname{Ch}(P+Q)=[\det(\alpha(k_{2}))]=[\det(\beta(k_{2}))\det(\gamma(k_{2}))]=[\det(\beta(k_{2}))]+[\det(\gamma(k_{2}))].\qed$
3. The topological invariant
In a milestone paper [kane2005z], Kane and Mele demonstrated that it is possible to associate a $\mathbb{Z}_{2}$-valued topological invariant to a specific discrete honeycomb model. The methods introduced in that work were subsequently extended to a broader class of models in several later papers, including [fu2007topological], [freed2013twisted], [graf2013bulk], [de2015classification], [fiorenza2016z], and [monaco2015symmetry], each providing a distinct formulation of the invariant and a different interpretation of the physical information it encodes. In this section, we will present yet another formulation of the invariant, starting from a continuous projection-valued map
$P:\mathbb{T}^{2}\to\operatorname{Proj}_{m}(\mathcal{H})$
subject to the symmetry constraint described in Definition 1.4. We also note, as Lemma A.1 guarantees, that $\dim(\operatorname{Ran}(P(0,0)))=m$ is always even; hence, from this point onward, we set $m=2n$.
Consider a time-reversal symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ and a unitary-valued map $U:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$—obtained, for example, by using Proposition 2.3—such that
$P(t,k_{2})U(t,k_{2})=U(t,k_{2})P(0,0)\quad\mbox{and}\quad TU(t,k_{2})=U(-t,-k_{2})T\quad\forall t\in[-\pi,\pi],k_{2}\in\mathbb{T}^{1}.$
Then we can use Lemma A.1 to fix a quaternionic basis of $\operatorname{Ran}(P(0,0))$, i.e., a basis compatible with the quaternionic structure induced by the time-reversal operator $T$: fix an orthonormal set $\{u_{j}\}_{j\in\{1,\cdots n\}}$, and define $u_{n+j}=Tu_{j}$ for $j\in\{1,\cdots,n\}$ to have a complete basis of $\operatorname{Ran}(P(0,0))$. Then we can get a continuous collection of orthonormal bases of $P(t,k_{2})$ by setting, for all $t\in[-\pi,\pi],k\in\mathbb{T}^{1},j\in\{1,\cdots,2n\}$
$u_{j}(t,k_{2})=U(t,k_{2})u_{j}\quad\mbox{with}\quad Tu_{j}(t,k_{2})=\sum_{a=1}^{2n}J_{j,a}u_{a}(-t,-k_{2})$
where $J$ is as in (8). The aperiodicity in $k_{1}$ is expressed by a continuous family of matching matrices $\alpha(k_{2})\in U(2n)$ such that
$u_{j}(-\pi,k_{2})=\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u_{a}(\pi,k_{2}).$
This family satisfies the symmetry constraint $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$ because:
$\begin{split}\sum_{a=1}^{2n}[J]_{j,a}u_{a}(\pi,-k_{2})&=Tu_{j}(-\pi,k_{2})=T\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u(\pi,k_{2})\\ &=\sum_{a,b=1}^{2n}\overline{[\alpha(k_{2})]_{j,a}}[J]_{a,b}u_{b}(-\pi,-k_{2})=\\ &=\sum_{a,c=1}^{2n}\left[\overline{\alpha(k_{2})}(J)\right]_{j,a}[\alpha(-k_{2})]_{a,c}u_{c}(\pi,-k_{2})=\\ &=\sum_{a=1}^{2n}\left[\overline{\alpha(k_{2})}(J)\alpha(-k_{2})\right]_{j,a}u_{a}(\pi,-k_{2})\end{split}$
Using this symmetry constraint we can find two matrices $\gamma_{0},\gamma_{\pi}$ such that $\alpha(0)=J\gamma_{0}^{t}J^{-1}\gamma_{0}$ and same for $\gamma_{\pi}$. For example, since $J\alpha(0)^{t}J^{-1}=\alpha(0)$, we can find a self-adjoint matrix $L$ such that $\alpha(0)=e^{iL}$ and $JL^{t}J^{-1}=L$, then $\gamma_{0}=e^{iL/2}$ will work. It is obvious that $\det(\alpha(0))=\det(\gamma_{0})^{2}$ and $\det(\alpha(\pi))=\det(\gamma_{\pi})^{2}$. If $\det(\gamma_{0})=e^{i\lambda_{0}}$ and $\det(\gamma_{\pi})=e^{i\lambda_{\pi}}$ for two real numbers $\lambda_{0},\lambda_{\pi}\in\mathbb{R}$, we can choose a continuous function $\mu:[0,\pi]\to\mathbb{R}$ such that $\det(\alpha(k_{2}))=e^{i\mu(k_{2})}$ and $\mu(0)=2\lambda_{0}$. Then it holds that $e^{i\mu(\pi)}=e^{i2\lambda_{\pi}}$, so $\mu(\pi)-2\lambda_{\pi}\in 2\pi\mathbb{Z}$.
Definition 3.1 ($\delta(P)$).
Under these hypotheses, we will define:
$\delta(P)=e^{i\frac{\left(2\lambda_{\pi}-\mu(\pi)\right)}{2}}\in\{\pm 1\}=\mathbb{Z}_{2}.$
Proposition 3.2.
The quantity defined previously does not depend on the choices of $\lambda_{0},\lambda_{\pi}$ $\gamma_{0},\gamma_{\pi}$.
Proof.
If we took $\lambda^{\prime}(0)$, $\lambda^{\prime}(\pi)$ such that $e^{i\lambda_{0}}=e^{i\lambda^{\prime}(0)}$ and $e^{i\lambda_{\pi}}=e^{i\lambda^{\prime}(\pi)}$, we would have that $\lambda(0)=\lambda^{\prime}(0)+2\pi l$, $\lambda_{\pi}=\lambda^{\prime}(\pi)+2\pi m$ for some $l,m\in\mathbb{Z}$. This means that we need to impose $\mu^{\prime}(k)=\mu(k)+4\pi$ and obtain:
$\delta^{\prime}=e^{i\frac{2\lambda^{\prime}(\pi)-\mu^{\prime}(\pi)}{2}}=e^{i\frac{2\lambda_{\pi}+4\pi m-\mu(\pi)-4\pi l}{2}}=\delta\cdot e^{i2\pi(m-l)}=\delta$
Instead if we took another $\gamma^{\prime}_{0}$ such that $\alpha(0)=J\gamma_{0}^{t}J^{-1}\gamma_{0}=J(\gamma^{\prime}_{0})^{t}J^{-1}\gamma^{\prime}_{0}$, then it is true that:
$(\gamma^{\prime}_{0})^{t}J\gamma^{\prime}_{0}=\gamma_{0}^{t}J\gamma_{0}\Leftrightarrow J=(\gamma_{0}(\gamma_{0}^{\prime})^{-1})^{t}J(\gamma_{0}(\gamma_{0}^{\prime})^{-1})$
This means that $\gamma_{0}(\gamma^{\prime}_{0})^{-1}\in\mathrm{Sp}(2n)$, so $\det(\gamma_{0}(\gamma^{\prime}_{0})^{-1})=1$ and therefore $\det(\gamma_{0})=\det(\gamma^{\prime}_{0})$. Since the same holds for $\gamma_{\pi}$, we have that $\det(\gamma_{0}),\det(\gamma_{\pi})$ depends only on $\alpha(0),\alpha(\pi)$. ∎
The invariant can also be formulated as the parity of a Chern number. In particular, if $P^{\pm}:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ is a pair of projection-valued maps such that $TP^{+}(k)=P^{-}(-k)T$ and with $P(k)=P^{+}(k)+P^{-}(k)$, then $\delta(P)=e^{i\pi\operatorname{Ch}(P^{-})}$. However, this formulation requires an entire section of the paper to be fully explored. Nonetheless, we can borrow the main result of the next section to prove the following Proposition.
Proposition 3.3.
If two time-reversal symmetric projection-valued maps $P_{0},P_{1}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ are unitarily equivalent or homotopic, then $\delta(P_{0})=\delta(P_{1})$.
Proof.
Suppose that there is a unitary-valued map $U:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ such that $P_{1}(k)=U(k)P_{0}(k)U(k)^{-1}$ with TU(k)=U(k)TTU(k)=U(-k)T. Then we can use Theorem 4.2 to decompose $P_{0}=P_{0}^{-}(k)\oplus P_{0}^{+}(k)$ with $P_{0}^{\pm}:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$ and $TP_{0}^{+}(k)=P_{0}^{-}(-k)T$. Clearly
$P_{1}(k)=U(k)P_{0}^{-}(k)U(k)^{-1}\oplus U(k)P_{0}^{+}(k)U(k)^{-1}$
is a decomposition of $P_{1}(k)$ such that $TP_{1}^{+}(k)=P_{1}^{-}(-k)T$. Since the Chern number is stable under unitary equivalences (Proposition 2.5), we have that $\operatorname{Ch}(P_{0}^{-})=\operatorname{Ch}(P_{1}^{-})$. So, it follows that:
$\delta(P_{1})=e^{i\pi\operatorname{Ch}(P_{1}^{-})=}e^{i\pi\operatorname{Ch}(P_{0}^{-})}=\delta(P_{0}).$
Instead, suppose that there is an homotopy $P_{t}(k)$ with $TP_{t}(k)=P_{t}(-k)T$ for $t\in[0,1]$ that connects $P_{0}$ with $P_{1}$. then we can divide $[0,1]$ into a partition $t_{j},t_{j+1}]_{j\in\{0,\cdots,J-1\}}$ such that $t_{0}=0,t_{J}=1$ and such that $\|P_{t}(k)-P_{t_{j}}(k)\|<1$ for all $t\in[t_{j},t_{j+1}],k\in\mathbb{T}^{2}$. Then we can use the Kato-Nagy formula 7 to define the unitary-valued maps
$U_{j}(k)=P_{t_{j+1}}(k)P_{t_{j}}(k)+(\mathds{1}-P_{t_{j+1}}(k))(\mathds{1}-P_{t_{j}}(k))\left[\mathds{1}-(P_{t_{j}}(k)-P_{t_{j+1}}(k))^{2}\right]^{-1/2}$
for all $j\in\{0,\cdots,J-1\}$. It is easy to see that $U_{j}(k)$ is a symmetric unitary equivalence between $P_{t_{j}}(k)$ and $P_{t_{j+1}}(k)$ for all $j\in\{0,\cdots,J-1\}$, so, using the previous point, we see that $\delta$ is preserved at each step and therefore $\delta(P_{0})=\delta(P_{1})$. ∎
4. Splitting of time-reversal symmetric projections
In this section, we want to describe how it is possible to express the invariant $\delta$ as the parity of a Chern number. To do so, we need to use a specific formulation of the Chern number and we also need to give an algorithm that allows us to decompose $P$ as a pair $P=P^{+}+P^{-}$.
Remark 4.1.
As stated in [panati2007triviality, monaco2015symmetry], if $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ is a time-reversal symmetric projection-valued map, then $\operatorname{Ch}(P)=0$.
Indeed, we can use Proposition 2.3 to define a unitary-valued map $U:[-\pi,\pi]\times\mathbb{T}^{1}\to\mathcal{U}(\mathcal{H})$ with $P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}$ and $TU(t,k_{2})=U(-t,-k_{2})T$. Then if $\{v_{j}\}_{j\in\{1,\cdots,n\}}$ is a quaternionic orthonormal basis of $\operatorname{Ran}(P(0,0))$ and $v_{j+n}=Tv_{j}$ for $j\in\{1,\cdots,n\}$ as in Lemma A.1, we will have the unitary matrix $\alpha\in U(2n)$ with:
$\begin{split}[\alpha]_{a,b}(k_{2})&=\left\langle v_{a},U(-\pi,k_{2})^{-1}U(\pi,k_{2})v_{b}\right\rangle=\left\langle U(-\pi,k_{2})v_{a},U(\pi,k_{2})v_{b}\right\rangle=\\ &=\left\langle TU(\pi,k_{2})v_{b},TU(-\pi,k_{2})v_{a}\right\rangle=\left\langle U(-\pi,-k_{2})v_{b},U(\pi,-k_{2})v_{a}\right\rangle=\\ &=[\alpha]_{b,a}(-k_{2}).\end{split}$
This means that $\det(\alpha(k_{2}))=\det(\alpha(-k_{2}))$, but a map with this property has a winding number equal to zero due to Lemma A.4.
Theorem 4.2.
Given a time-reversal symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ and an integer $h\in\mathbb{Z}$, if $e^{i\pi h}=\delta(P)$, it is always possible to decompose $P$ as $P(k)=P^{+}(k)\oplus P^{-}(k)$ with $\operatorname{Ch}(P^{-})=h$. Vice versa, if $P$ can be decomposed as $P(k)=P^{+}(k)\oplus P^{-}(k)$, then $e^{i\pi\operatorname{Ch}(P^{-})}=\delta(P)$.
Proof.
We want to start using the same notation used in Definition 3.1, which means that $P(t,k_{2})$ is spanned by $\{u_{j}(t,k_{2})\}_{j\in\{1,\cdots,2n\}}$ with constraints:
$Tu_{j}(t,k_{2})=\sum_{a=1}^{2n}[-J]_{j,a}u_{a}(-t,-k_{2}),\quad u_{j}(-\pi,k_{2})=\sum_{a=1}^{2n}[\alpha(k_{2})]_{j,a}u_{a}(\pi,k_{2})$
where $\alpha$ is a periodic unitary-valued map such that $J=\overline{\alpha(k_{2})}J\alpha(-k_{2})$ and $J=\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}$.
Equivalent conditions.  Anticipating some results from the next section, we notice that the existence of a decomposition is tantamount to the existence of a symmetric frame with appropriate pseudo-periodic conditions. In fact, given a decomposition $P(k)=P^{+}(k)\oplus P^{-}(k)$, we can apply Theorem 5.1 to $P(k_{1},k_{2})$ to construct a frame $\{v_{j}(k_{1},k_{2})\}_{j\in\{1,\cdots,n\}}$ of $P^{-}(k_{1},k_{2})$ with complete periodicity except for the first vector, that instead satisfies
$v_{1}(\pi,k_{2})=e^{i\operatorname{Ch}(P^{-})k_{2}}v_{1}(-\pi,k_{2})=e^{ihk_{2}}v_{1}(-\pi,k_{2}).$
Then
$\{v_{n+j}(t,k_{2})=-Tv_{j}(-t,-k_{2})\}_{j\in\{1,\cdots,n\}}$
is a frame of $P^{+}(t,k_{2})=T^{-1}P^{-}(-t,-k_{2})T$ with complete periodicity except for the first vector that satisfies
$v_{n+1}(\pi,k_{2})=e^{-i\operatorname{Ch}(P^{-})k_{2}}v_{n+1}(-\pi,k_{2})=e^{-ihk_{2}}v_{n+1}(-\pi,k_{2})\,.$
The union of the two collections of vectors will constitute a symmetric frame for $P$ and this also implies that $\operatorname{Ch}(P^{+})=-\operatorname{Ch}(P^{-})$. Vice versa, if we have a frame with these properties, we can take $P^{-}(k)$ as the span of the first $n$ vectors and $P^{+}(k)$ as the span of the remaining $n$ vectors.
At the same time, the existence of a frame of this type is tantamount to the existence of a unitary-valued map $\beta:[-\pi,\pi]\times\mathbb{T}^{1}\to U(2n)$ such that $v_{j}(t,k_{2})=\sum_{a=1}^{2n}[\beta(t,k_{2})]_{j,a}u_{a}(t,k_{2})$. The frame is symmetric if and only if $J\beta(t,k_{2})=\overline{\beta(-t,-k_{2})}J$ because:
$\begin{split}Tv_{j}(t,k_{2})&=\sum_{a=1}^{2n}[J]_{j,a}v_{a}(-t,-k_{2})\Leftrightarrow\\ \Leftrightarrow\sum_{a=1}^{2n}\overline{[\beta(t,k_{2})]_{j,a}}Tu_{a}(t,k_{2})&=\sum_{a,b=1}^{2n}[J]_{j,a}[\beta(-t,-k_{2})]_{a,b}u_{b}(-t,-k_{2})\Leftrightarrow\\ \Leftrightarrow\sum_{b=1}^{2n}\left[\overline{\beta(t,k_{2})}(J)\right]_{j,b}u_{b}(-t,-k_{2})&=\sum_{b=1}^{2n}\left[J\beta(-t,-k_{2})\right]_{j,b}u_{b}(-t,-k_{2})\end{split}$
Instead, the frame satisfies the pseudoperiodic condition if and only if
$\beta(\pi,k_{2})=\Lambda(k_{2})\beta(-\pi,k_{2})\alpha(k_{2})$
where $\Lambda(k)$ is a diagonal matrix with
$\Lambda(k)=\operatorname{diag}(e^{ihk_{2}},\underbrace{1,\cdots,1}_{n-1},e^{-ihk_{2}},\underbrace{1,\cdots,1}_{n-1}).$
This is true because:
$\begin{split}v_{j}(\pi,k_{2})&=\sum_{a=1}^{2n}[\Lambda(k_{2})]_{j,a}v_{a}(-\pi,k_{2})\Leftrightarrow\\ \sum_{a=1}^{2n}[\beta(\pi,k_{2})]_{j,a}u_{a}(\pi,k_{2})&=\sum_{a=1}^{2n}[\Lambda(k_{2})\beta(-\pi,k_{2})]_{j,a}u_{a}(-\pi,k_{2})\Leftrightarrow\\ \sum_{a=1}^{2n}[\beta(\pi,k_{2})]_{j,a}u_{a}(\pi,k_{2})&=\sum_{a=1}^{2n}[\Lambda(k_{2})\beta(-\pi,k_{2})\alpha(k_{2})]_{j,a}u_{a}(\pi,k_{2})\end{split}$
However, the symmetry condition forces a constraint on $\beta$: in fact if $J\beta(0,k_{2})=\overline{\beta(0,-k_{2})}J$, then $\beta(0,0),\beta(0,\pi)\in\mathrm{Sp}(2n)$. This implies that $\det(\beta(0,0))=\det(\beta(0,\pi))=1$. Moreover, $\det(\beta(0,k_{2}))=\overline{\det(\beta(0,-k_{2})}$, so we can use the Lemma A.3 to conclude that the winding number in the second direction $\det(\beta(t,\cdot))]$ must be even.
Decomposition when $\delta(P)=1$. Now we want to investigate under which assumption it is possible to find such a $\beta$ when $h$ is even. First of all, we can look at the points $(\pi,0),(\pi,\pi)$ in the torus. There it must be true that:
$\beta(\pi,0)=\Lambda(0)\beta(-\pi,0)\alpha(0),\quad\beta(\pi,\pi)=\Lambda(\pi)\beta(-\pi,\pi)\alpha(\pi).$
However, if $h\in 2\mathbb{Z}$, then $\Lambda(0)=\Lambda(\pi)=\mathbb{I}_{2n}$. If we use the periodicity in $k_{2}$ and the symmetry constraint, we get that $\beta$ must satisfy
$\beta(\pi,k_{2})=\Lambda(k_{2})J\overline{\beta(\pi,-k_{2})}J^{-1}\alpha(k_{2})$
and in particular
$J\beta(\pi,0)^{t}J^{-1}\beta(\pi,0)=\alpha(0),\quad J\beta(\pi,\pi)^{t}J^{-1}\beta(\pi,\pi)=\alpha(\pi).$
This means that, since $\det(\Lambda(k_{2}))\equiv 1$, it is true that
$\begin{cases}\det(\beta(\pi,k_{2}))=\overline{\det(\beta(\pi,-k_{2}))}\det(\alpha(k_{2}))\\ \det(\beta(\pi,0))^{2}=\det(\alpha(0))\\ \det(\beta(\pi,\pi))^{2}=\det(\alpha(\pi))\end{cases}$
Now we can choose a function $\lambda:[0,\pi]\to\mathbb{R}$ such that $\det(\beta(\pi,k_{2}))=e^{i\lambda(k_{2})}$, as well as a function $\mu:[-\pi,\pi]\to\mathbb{R}$ such that $\det(\alpha(k_{2}))=e^{i\mu(k_{2})}$ and with $\mu(0)=2\lambda(0)$. This means that:
$\det(\beta(\pi,k_{2}))=\begin{cases}e^{i\lambda(k_{2})}&\mbox{for }k_{2}\in[0,\pi]\\ e^{-i\lambda(-k_{2})+i\mu(-k_{2})}&\mbox{for }k_{2}\in[-\pi,0]\end{cases}$
So we can write $\det(\beta(\pi,k_{2}))=e^{i\phi(k_{2})}$ where $\phi:[-\pi,\pi]\to\mathbb{R}$ is a continuous but not periodic function such that:
$\phi(k_{2})=\begin{cases}\lambda(k_{2})&\mbox{for }k_{2}\in[0,\pi]\\ \lambda(-k_{2})+\mu(-k_{2})&\mbox{for }k_{2}\in[-\pi,0]\end{cases}$
This implies that
$\det(\beta(\pi,\cdot))]=\frac{\lambda(\pi)-[-\lambda(\pi)+\mu(\pi)]}{2\pi}=\frac{2\lambda(\pi)-\mu(\pi)}{2\pi}$
So, if we interpret $\beta(\pi,0)$ as $\gamma_{0}$ and $\beta(\pi,\pi)$ as $\gamma_{\pi}$, then we can say that $\lambda(0)=\lambda_{0},\lambda(\pi)=\lambda_{\pi}$ and obtain that
$e^{i\pi[\det(\pi,k_{2})]}=e^{i\pi\frac{2\lambda(\pi)-\mu(\pi)}{2\pi}}=\delta(P)$
This means that if $\delta(P)$ is odd, it is impossible to find $\beta$ because $\delta$ does not depend on the choices of $\beta(\pi,0)$ and $\beta(\pi,\pi)$ (Proposition 3.2) and the symmetry constraint imposed that $\det(\beta(t,\cdot))]$ must be even for every $t$. Instead if $\delta(P)=2h$ is even we can use the Sylvester theorem to find two self-adjoint operators $L,R$ with $JLJ^{-1}=L^{t},JRJ^{-1}=R^{t}$ and such that $\alpha(0)=e^{iL},\alpha(\pi)=e^{iR}$, and then define:
$\beta(\pi,k_{2})=\begin{cases}\exp\left[i\frac{\pi-k_{2}}{2\pi}L+i\frac{k_{2}}{2\pi}R\right]&\mbox{for }k_{2}\in[0,\pi]\\ \Lambda(k_{2})J\overline{\beta(\pi,-k_{2})}J^{-1}\alpha(k_{2})&\mbox{for }k_{2}\in[-\pi,0]\end{cases}$
This is continuous and periodic in $k_{2}$ because when $k_{2}=0$ we get:
$e^{iL/2}=\Lambda(0)Je^{-i\overline{L}/2}J^{-1}e^{iL}=e^{-iJL^{t}J^{-1}/2}e^{iL}=e^{iL/2}$
and similar computations hold for $k_{2}=\pi$. The previous argument states that $\det(\beta(\pi,k_{2}))]$ is even equal to $2r$. Then we can also define
$\beta(0,k_{2})=\operatorname{diag}(e^{irk_{2}},\underbrace{1,\cdots,1}_{n-1},e^{irk_{2}},\underbrace{1,\cdots,1}_{n-1})$
and it is trivial to check that $\beta(0,k_{2})$ satisfies the symmetry constraint and has $\det(\beta(0,k_{2}))]=2r=[\det(\beta(\pi,k_{2}))]$. Finally we can use Theorem A.7 to find a homotopy $\beta_{t}(k_{2})$ in $U(2n)$ with $\beta_{0}(k_{2})=\beta(0,k_{2}),\beta_{1}(k_{2})=\beta(\pi,k_{2})$ and then define:
$\beta(t,k_{2})=\begin{cases}\beta_{t/\pi}(k_{2})&\mbox{for }t\in[0,\pi]\\ J^{-1}\overline{\beta_{-t/\pi}(-k_{2})}J&\mbox{for }t\in[-\pi,0]\end{cases}$
This satisfies all the desired conditions, so if $\delta(P)$ is even, we can build a decomposition with $\operatorname{Ch}(P^{-})=h$ for any even number $h$.
Decomposition when $\delta(P)=-1$.  Now we want to investigate under which assumption it is possible to find a decomposition with $\operatorname{Ch}(P^{-})$ odd. It is immediate to check that the conditions are identical to the even case with a small but crucial difference. If $\operatorname{Ch}(P^{-})$ is even, it holds that:
$\Lambda(0)=\mathbb{I}_{2n},\quad\Lambda(\pi)=\begin{pmatrix}-1&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&-1&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}$
Once again, if we use the periodicity in $k_{2}$ and the symmetry constraint, we get that $\beta$ must satisfy
$\beta(\pi,k_{2})=\Lambda(k_{2})J\overline{\beta(\pi,-k_{2})}J^{-1}\alpha(k_{2})$
and in particular
$J\beta(\pi,0)^{t}J^{-1}\beta(\pi,0)=\alpha(0),\quad J\beta(\pi,\pi)^{t}J^{-1}\Lambda(\pi)\beta(\pi,\pi)=\alpha(\pi).$
With some computation, it is easy to check that
$\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}\begin{pmatrix}-1&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&-1&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}$=
$\begin{pmatrix}i&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&i&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}^{t}\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}\begin{pmatrix}i&0&0&0\\ 0&\mathbb{I}_{n-1}&0&0\\ 0&0&i&0\\ 0&0&0&\mathbb{I}_{n-1}\end{pmatrix}$
meaning that if
$Y=\operatorname{diag}(i,\underbrace{1,\cdots,1}_{n-1},i,\underbrace{1,\cdots,1}_{n-1})$
we have that $\det(Y)=-1$ and the last condition can be rewritten as:
$J(Y\beta(\pi,\pi))^{t}J^{-1}(Y\beta(\pi,\pi))=\alpha(\pi).$
We can replicate the previous step of the proof to show that it is possible to define $\beta$ if and only if $\frac{2\lambda(\pi)-\mu(\pi)}{2\pi}\in 2\mathbb{Z}.$ In this case we can interpret $\beta(\pi,0)$ as $\gamma_{0}$ and $Y\beta(\pi,\pi)$ as $\gamma_{\pi}$. Then we can impose $\lambda_{0}=-i\ln(\det(\pi,0))=\lambda(0)$, so that $\mu$ does not need to be changed, and $\lambda_{\pi}=-i\ln(\det(Y\beta(\pi,\pi)))=\lambda(\pi)+\pi$ because
$\det(Y\beta(\pi,\pi))=-\det(\beta(\pi,\pi))=e^{i(\lambda(\pi)+\pi)}.$
We conclude that, in this case, we have that:
$\exp\left(i\frac{2\lambda(\pi)-\mu(\pi)}{2}\right)=\exp\left(i\frac{2\lambda_{\pi}-\mu(\pi)-\pi}{2}\right)=-\delta(P)$
Since we need an even winding number of $\det(\beta(\pi,k_{2}))]$, if $\delta(P)=-1$ it is impossible to construct a frame with $h$ odd, instead if $\delta(P)=1$ we can proceed exactly as before and obtain the thesis. Namely, it is possible to decompose $P=P^{+}\oplus P^{-}$ with $\operatorname{Ch}(P^{-})$ odd if and only if $\delta(P)$ is odd. ∎
Lemma 4.3.
Consider two time-reversal symmetric projection-valued maps $P,Q:\mathbb{T}^{2}\mapsto\operatorname{Proj}_{2n}(\mathcal{H})$ such that $P(k)Q(k)\equiv 0$: then $\delta(P)\delta(Q)=\delta(P+Q)$.
Proof.
We can use the previous theorem to find two decompositions $P(k)=P^{+}(k)\oplus P^{-}(k)$, $Q(k)=Q^{+}(k)\oplus Q^{-}(k)$. Then it is obvious that we can find a decomposition of $P(k)\oplus Q(k)$ as:
$P(k)\oplus Q(k)=[P^{-}(k)\oplus Q^{-}(k)]\oplus[P^{+}(k)\oplus Q^{+}(k)]$
Now we can use Lemma 2.6, namely $\operatorname{Ch}(P^{-}\oplus Q^{-})=\operatorname{Ch}(P^{-})+\operatorname{Ch}(Q^{-})$. So we can conclude using the previous theorem:
$\delta(P\oplus Q)=e^{i\pi\operatorname{Ch}(P^{-}\oplus Q^{-})}=e^{i\pi\operatorname{Ch}(P^{-})}\cdot e^{i\pi\operatorname{Ch}(Q^{-})}=\delta(P)\cdot\delta(Q).\qed$
5. Pseudo-periodic frames
One approach to the topological study of projection-valued maps is to investigate under which assumption is it possible to build a periodic and symmetric frame for a given projection-valued map. To be precise, given $P:\mathbb{T}^{d}\to\operatorname{Proj}_{n}(\mathcal{H})$, the goal is to find a periodic and continuous (or analytic) collection of orthonormal vectors $\{v_{j}(k)\}_{j\in 1,\cdots,n}$ that span $\operatorname{Ran}(P(k))$ for all $k\in\mathbb{T}^{d}$. In some circumstances, this operation turns out to be impossible, and non trivial topological invariants like $\operatorname{Ch}$ and $\delta$ are seen as topological obstructions in the construction of a frame [panati2007triviality, monaco2015symmetry, fiorenza2016z].
More recently, a more refined goal has been researched, namely how close one can get to a global, symmetric, continuous, and periodic frame when its construction is topologically obstructed: see e.g. [cornean2019parseval, auckley2018parseval]. One relaxes the notion of periodicity to one of pseudo-periodicity (or periodicity up to a phase) for the minimal number possible of vectors in the basis. For example, for a non-symmetric projection-valued map the following result from [monaco2023topology] holds.
Theorem 5.1 (non-symmetric pseudo-periodic frame).
Consider a projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{n}(\mathcal{H})$. Then it is always possible to construct a pseudo-periodic frame $\{v_{j}(k)\}_{j\in\{1,\cdots,n\}}$ such that all vectors are periodic in $k$ except the first that satisfies
$v_{1}(\pi,k_{2})=e^{i\operatorname{Ch}(P)k_{2}}v_{1}(-\pi,k_{2}),\quad v_{1}(k_{1},0)=v_{1}(k_{1},2\pi)\quad\forall k_{1},k_{2}\in\mathbb{T}^{2}.$
Proof.
Using Proposition 2.3 we can build a unitary-valued map periodic in $k_{2}$ such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}\quad\mbox{for}\quad(t,k_{2})\in[-\pi,\pi]\times\mathbb{T}^{1}.$
If $\{v_{j}\}_{j\in\{1,\cdots,n\}}$ is an orthonormal basis of $P(0,0)$, then $\{u_{j}(t,k_{2})=U(t,k_{2})v_{j}\}_{j\in\{1,\cdots,n\}}$ constitutes an orthonormal basis of $P(t,k_{2})$ that depends continuously in $t$ and $k_{2}$. This is periodic in $k_{2}$ but in general will not be periodic in $t$. The aperiodicity is once again measured by the matching matrix $\alpha(k_{2})\in U(n)$, with:
$u_{j}(\pi,k_{2})=\sum_{a=1}^{n}\alpha_{j,a}(k_{2})u_{j}(-\pi,k_{2}).$
Since $\det(\alpha(k_{2}))]=\operatorname{Ch}(P)$ the matrix:
$\beta(k_{2})=\begin{pmatrix}e^{(i\operatorname{Ch}(P)k_{2})}&0\\ 0&\mathds{1}_{n-1}\end{pmatrix}\alpha(k_{2})^{-1}$
is such that $\det(\beta(k_{2}))]=0$. So, Using Theorem A.7, there is a homotopy $\beta_{s}(k_{2})$ with $\beta_{1}(k_{2})=\beta(k_{2})$ and $\beta_{0}(k_{2})=\mathds{1}$. Therefore, the vectors
$v_{j}(t,k_{2})=\sum_{a=1}^{n}\left[\beta_{(t+\pi)/2\pi}(k_{2})\right]_{j,a}u_{j}(t,k_{2})\quad\mbox{for}\quad j\in\{1,\cdots,n\}$
constitute an orthonormal basis of $\operatorname{Ran}(P(t,k_{2}))$ and are such that:
$v_{j}(\pi,k_{2})$ $=\sum_{a=1}^{n}[\beta(k_{2})]_{j,a}u_{a}(\pi,k_{2})=\sum_{a,b=1}^{n}[\beta(k_{2})]_{j,a}\alpha_{a,b}(k_{2})u_{b}(-\pi,k_{2})=$
$=\sum_{a=1}^{n}[\beta(k_{2})\alpha(k_{2})]_{j,a}u_{a}(-\pi,k_{2})=\begin{cases}e^{i\operatorname{Ch}(P)k_{2}}v_{1}(-\pi,k_{2})&\mbox{for }j=1,\\ v_{j}(-\pi,k_{2})&\mbox{otherwise}.\end{cases}\qed$
Instead, in presence of a time-reversal symmetry $T$ such that $T^{2}=-\mathds{1}$, the most natural symmetric constraint we can impose is that the frame must be composed by pairwise symmetric vectors $v(k),u(k)$ such that Tv(k)=u(k)Tv(k)=u(-k). This is the optimal condition because it is not possible to relate a vector to itself by imposing Tv(k)=v(k)Tv(k)=v(-k). In fact, it would lead to $Tv(0)=v(0)$ and consequently to $v(0)=T^{2}v(0)=v(0)$ which is absurd. This relation can instead be imposed when $T^{2}=\mathds{1}$, as discussed in [cornean2016construction] and [fiorenza2016z].
Theorem 5.2 (Pseudo-periodic symmetric frame).
Consider a time-reversal symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$. Then, it is always possible to construct a symmetric and pseudo-periodic frame
$\{v_{1},\cdots,v_{n}(k),-Tv_{1}(-k),\cdots,-Tv_{n}(-k)\}$
of $P(k)$ such that, if $\delta(P)=1$, the frame has full periodicity, while if $\delta(P)=-1$, the frame is periodic with the only exceptions being:
$v_{1}(\pi,k_{2})=e^{ik_{2}}v_{1}(-\pi,k_{2})\quad\text{and}\quad v_{n+1}(\pi,k_{2})=e^{-ik_{2}}v_{n+1}(-\pi,k_{2})\quad\forall k_{2}\in\mathbb{T}^{1}.$
Proof.
The construction of the symmetric frame can be done by applying Theorem 4.2: if $\delta(P)=1$ we can build a decomposition with any even Chern number, in particular with a vanishing one. Instead, if $\delta(P)=-1$ we can build a decomposition $P=P^{-}\oplus P^{+}$ with $P^{-}$ having any odd Chern number, in particular with $\operatorname{Ch}(P^{-})=1$. To conclude, we need to apply Theorem 5.1 to $P^{-}$ and then symmetrize the frame obtained. ∎
Remark 5.3 (Symmetry vs periodicity).
Theorems 5.1 and 5.2 together with Remark 4.1 yield an interesting fact. If we have a symmetric projection-valued map $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ we have that $\operatorname{Ch}(P)=0$. So, it is possible to construct a periodic and continuous frame. However, if we are looking for a frame that respects the symmetry $T$, it is not possible to construct a continuous, symmetric, and periodic frame when $\delta(P)=-1$. This means that there is a natural question that arises: what is the natural conflict between symmetry and periodicity? The answer follows naturally from Theorem 5.2. In fact, if we look at the non-periodic vectors $v_{1},v_{n+1}$ we have that they span a time-reversal symmetric projection-valued map $\tilde{P}:\mathbb{T}^{2}\to\operatorname{Proj}_{2}(\mathcal{H})$ such that $\delta(\tilde{(}P))=-1$ and that contains the non-trivial topological information of $P$. Instead, the other periodic vectors $v_{2},\cdots,v_{n-1},v_{n+2},\cdots,v_{2n}$ span a time-reversal symmetric projection-valued map $\hat{P}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n-2}(\mathcal{H})$ such that $\delta(\hat{P})=1$. So if one looks for periodicity over symmetry it is possible to construct a frame in which only two vectors are not pairwise symmetric while the others are perfectly continuous, symmetric and periodic.
Following this lines, by repeating the construction of [cornean2019parseval], one can use the same “Hilbert space doubling” trick employed there to the map $\tilde{P}$ in order to span it with four periodic, non-orthogonal vectors, which however come in time-reversal-symmetry-related Kramers pair. Together with the symmetric and periodic frame for $\hat{P}$, these $m+2$ vector constitute a continuous, symmetric and periodic Parseval frame spanning the projection $P$.
6. Homotopy problem
In Definition 1.4 we gave a notion of symmetric unitary equivalence and a notion of symmetric homotopy. The goal of this section is to prove that $\delta$ is the only topological information needed in order to connect two different projection-valued maps via symmetric homotopies or symmetric unitary equivalences. Before proving this result, we need a technical lemma that essentially allows us to construct symmetric, continuous, and periodic frames for infinite-dimensional time-reversal symmetric projection-valued maps.
Lemma 6.1.
If $\mathcal{H}$ is a separable Hilbert space with $\dim(\mathcal{H})=\infty$ and $P:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ is a continuous time-reversal symmetric projection-valued map with TP(k)=P(k)TTP(k)=P(-k)T, then there is always a continuous and periodic collection of orthonormal vectors $\{v_{j}(k)\}_{j\in\mathbb{N}}$ such that
$\{v_{1}(k),-Tv_{1}(-k),\cdots,v_{j}(k),-Tv_{j}(-k),\cdots\}$
is an orthonormal basis of $\ker(P(k))$.
Proof.
We can start by using Lemma A.1 to select a discrete quaternionic basis $\{v_{j}\}_{j\in\mathbb{N}}$ of $\ker(P(0,0))$. This basis, together with $\{Tv_{j}\}_{j\in\mathbb{N}}$, is a complex basis of $\ker(P(0,0))$. Then we can use Proposition 2.3 to construct a unitary-valued map $U(t,k_{2})$ such that
$P(t,k_{2})=U(t,k_{2})P(0,0)U(t,k_{2})^{-1}\quad\mbox{and}\quad TU(t,k_{2})=U(-t,-k_{2})T.$
We can construct an orthonormal basis of $\ker(P(t,k_{2}))$ using the vectors
$\{U(t,k_{2})v_{1},U(t,k_{2})Tv_{1},\cdots,U(t,k_{2})v_{j},U(t,k_{2})Tv_{j},\cdots\}.$
This is periodic in the second argument, but not in the first. To solve this problem, we notice that this family defines a family of natural isomorphisms:
$\begin{matrix}I(t,k_{2}):&\ker(P(t,k_{2}))&\to&l^{2}(\mathbb{N})\times l^{2}(\mathbb{N})\\ &U(t,k_{2})v_{j}&\mapsto&(\delta_{j},0)\\ &U(t,k_{2})Tv_{j}&\mapsto&(0,\delta_{j})\end{matrix}\quad\mbox{where }\delta_{j}(n)=\begin{cases}1\quad\mbox{if }n=j,\\ 0\quad\mbox{otherwise.}\end{cases}$
This family enjoys the following properties:
$TI(t,k_{2})^{-1}(\delta_{j},0)=TU(t,k_{2})v_{j}(0)=U(-t,-k_{2})Tv_{j}(0)=I(-t,-k_{2})^{-1}(0,\delta_{j})$
$TI(t,k_{2})^{-1}(0,\delta_{j})=TU(t,k_{2})Tv_{j}=-U(-t,-k_{2})v_{j}=I(-t,-k_{2})^{-1}(-\delta_{j},0)$
so by anti-linearity
$TI(t,k_{2})^{-1}=I(-t,-k_{2})^{-1}\mathcal{K}J\quad\mbox{where }J=\begin{pmatrix}0&\mathds{1}_{l^{2}(\mathbb{N})}\\ -\mathds{1}_{l^{2}(\mathbb{N})}&0\end{pmatrix}\quad\forall t\in[-\pi,\pi]$
and $\mathcal{K}$ is the standard complex conjugation. The aperiodicity is expressed by the fact that
$I(-\pi,k_{2})^{-1}I(\pi,k_{2})\neq\mathds{1}_{\mathcal{H}}.$
Using the spectral theorem, we know that there are two self-adjoint operators $L,R$ acting over $l^{2}(\mathbb{N})^{2}$ with
$I(\pi,0)I(-\pi,0)^{-1}=e^{iL}\quad\mbox{and}\quad I(\pi,\pi)I(-\pi,\pi)^{-1}=e^{iR}.$
Moreover, it is true that:
$\begin{split}\mathcal{K}J^{-1}I(\pi,0)I(-\pi,0)^{-1}&=I(-\pi,0)T^{-1}I(-\pi,0)^{1}=I(-\pi,0)I(\pi,0)^{-1}\mathcal{K}J^{-1}=\\ &=(I(\pi,0)I(-\pi,0)^{-1})^{-1}\mathcal{K}J^{-1}.\end{split}$
This means that $\mathcal{K}J^{-1}e^{iL}\mathcal{K}J=e^{-iL}$, so $e^{-iJ^{-1}\overline{L}J}=e^{-iL}$. Therefore, $L$ can be taken such that $J^{-1}\overline{L}J=L$. We can do the same for $R$ to have that $J^{-1}\overline{R}J=R$. Using them, we can define the unitary operator $\beta(k_{2})\in\mathcal{U}(l^{2}(\mathbb{N})^{2})$ with:
$\beta(k_{2})=\begin{cases}\exp\{\frac{i}{\pi}[(\pi-k_{2})\frac{L}{2}+k_{2}\frac{R}{2}]\}\quad&\mbox{if}\quad k_{2}\in[0,\pi]\\ \mathcal{K}J^{-1}I(-\pi,-k_{2})I(\pi,-k_{2})^{-1}\beta(-k_{2})\mathcal{K}J\quad&\mbox{if}\quad k_{2}\in[-\pi,0]\end{cases}$
and it is easily shown that $\beta(k_{2})$ is continuous and periodic in $k_{2}$ by applying the symmetry relations. Now, thanks to Theorem A.8 we know that there is a homotopy $\beta_{t}(k_{2})$ in $\mathcal{U}(l^{2}(\mathbb{N})^{2})$ with $\beta_{0}(k_{2})\equiv\mathds{1}_{l^{2}(\mathbb{N})^{2}}$ and $\beta_{1}(k_{2})=\beta(k_{2})$. To conclude the proof, we can define $\tilde{I}^{-1}(t,k_{2}):l^{2}(\mathbb{N})^{2}\to\ker(P(t,k_{2}))$ as:
$\tilde{I}(t,k_{2})^{-1}=\begin{cases}I(t,k_{2})^{-1}\beta_{t/\pi}(k_{2})&\mbox{for }t\in[0,\pi],k_{2}\in\mathbb{T}^{1}\\ I(t,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{-t/\pi}(-k_{2})\mathcal{K}J&\mbox{for }t\in[-\pi,0],k_{2}\in\mathbb{T}^{1}\end{cases}$
This $\tilde{I}$ is continuous for $t=0$ because $\beta_{0}(k_{2})\equiv\mathds{1}$, and is also periodic because, for $k_{2}\in[0,\pi]$, we have that
$\begin{split}\tilde{I}(-\pi,k_{2})&=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{1}(-k_{2})\mathcal{K}J=\\ &=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\mathcal{K}J^{-1}I(-\pi,k_{2})I(\pi,k_{2})^{-1}\beta(k_{2})\mathcal{K}J\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}\beta(k_{2})=\tilde{I}(\pi,k_{2})\end{split}$
while for $k_{2}\in[-\pi,0]$ we have that
$\begin{split}\tilde{I}(-\pi,k_{2})&=I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta_{1}(-k_{2})\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}I(\pi,k_{2})I(-\pi,k_{2})^{-1}\mathcal{K}J^{-1}\beta(-k_{2})\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}\mathcal{K}J^{-1}I(-\pi,-k_{2})I(\pi,-k_{2})^{-1}\beta(-k_{2})\mathcal{K}J=\\ &=I(\pi,k_{2})^{-1}\beta(k_{2})=\tilde{I}(\pi,k_{2})\end{split}$
Moreover, $\tilde{I}$ also satisfies a symmetry constraint:
$\begin{split}T\tilde{I}(t,k_{2})^{-1}&=\begin{cases}I(-t,-k_{2})^{-1}\mathcal{K}J\beta_{t/\pi}(k_{2})&\mbox{for }t\in[0,\pi]\\ I(-t,-k_{2})^{-1}\mathcal{K}J\mathcal{K}J^{-1}\beta_{-t/\pi}(-k_{2})\mathcal{K}J&\mbox{for }t\in[-\pi,0]\end{cases}\\ &=\begin{cases}I(-t,-k_{2})^{-1}\mathcal{K}J^{-1}\beta_{t/\pi}(k_{2})\mathcal{K}J\mathcal{K}J&\mbox{for }t\in[0,\pi]\\ I(-t,-k_{2})^{-1}\beta_{-t/\pi}(-k_{2})\mathcal{K}J&\mbox{for }t\in[-\pi,0]\end{cases}\\ &=\tilde{I}(-t,-k_{2})^{-1}\mathcal{K}J\end{split}$
Then $\{\tilde{v}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(\delta_{j},0)\}_{j\in\mathbb{N}}$ is the collection of vectors we were looking for, as it is periodic and continuous because so is $\tilde{I}$. Moreover, elements such as $(0,\delta_{j})$ together with $(\delta_{j},0)$ form an orthonormal basis of $l^{2}(\mathbb{N})^{2}$, so their images $\{\tilde{v}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(\delta_{j},0)\}_{j\in\mathbb{N}}$ together with $\{\tilde{u}_{j}(t,k_{2})=\tilde{I}(t,k_{2})^{-1}(0,\delta_{j})\}_{j\in\mathbb{N}}$ form a discrete orthonormal basis of $\ker(P(t,k_{2}))$. Most importantly, the relation $(\delta_{j},0)=\mathcal{K}J(0,\delta_{j})$ implies that $\tilde{v}_{j}(t,k_{2})=T\tilde{u}_{j}(-t,-k_{2})$ for all $j\in\mathbb{N},t\in[-\pi,\pi],k_{2}\in\mathbb{T}^{1}$. So, the vectors
$\{\tilde{v}_{1}(t),\tilde{u}_{1}(t),\cdots,\tilde{v}_{j}(t),\tilde{u}_{j}(t),\cdots\}$
are exactly the collection of orthonormal vectors we needed to complete the proof. ∎
Theorem 6.2.
Two time-reversal symmetric projection-valued maps $P_{0},P_{1}:\mathbb{T}^{2}\to\operatorname{Proj}_{2n}(\mathcal{H})$ are symmetrically unitarily equivalent and homotopic if and only if $\delta(P_{0})=\delta(P_{1})$.
Proof.
We already know one part of the statement: in fact, Proposition 3.3 states that if two maps are unitarily equivalent or homotopic, then their $\delta$ invariants are equal. So, we want to construct a symmetric unitary equivalence between $P_{0}$ and $P_{1}$ under the assumption that $\delta(P_{0})=\delta(P_{1})$. We can start by applying Theorem 5.2 to construct two (pseudo-periodic) frames
$\{v_{1}^{0}(k),\cdots,v_{n}^{0}(k),-Tv_{1}^{0}(-k),\cdots,-Tv_{n}^{0}(-k)\}$
$\{v_{1}^{1}(k),\cdots,v_{n}^{1}(k),-Tv_{1}^{1}(-k),\cdots,-Tv_{n}^{1}(-k)\}$
of $P_{0}(k)$ and $P_{1}(k)$ respectively, such that
$\begin{matrix}v_{1}^{0}(\pi,k_{2})=e^{i\pi h(\delta)k_{2}}v_{1}^{0}(-\pi,k_{2})\\ v_{1}^{1}(\pi,k_{2})=e^{i\pi h(\delta)k_{2}}v_{1}^{1}(-\pi,k_{2})\end{matrix}\quad\mbox{where }h(\delta)=\begin{cases}0&\mbox{if }\delta(P_{0})=\delta(P_{1})=1\\ 1&\mbox{if }\delta(P_{0})=\delta(P_{1})=-1\end{cases}$
Now we need to look at $\mathds{1}-P_{0}(k)$ and $\mathds{1}-P_{1}(k)$. These are once again two time-reversal symmetric projection-valued maps. If $\dim(\mathcal{H})<\infty$, thanks to lemma 4.3, it holds that
$\delta(\mathds{1}-P_{0})=\delta(\mathds{1})\cdot\delta(P_{0})=\delta(P_{0})=\delta(P_{1})=\delta(\mathds{1})\cdot\delta(P_{1})=\delta(\mathds{1}-P_{1}).$
So we can apply Theorem 5.2 to extend the previous frames obtaining:
$\{v_{1}^{0}(k),\cdots,-Tv_{n}^{0}(-k),v_{n+1}^{0}(k),\cdots,v_{\dim(\mathcal{H})/2}^{0}(k),-Tv_{n+1}^{0}(-k),\cdots-Tv_{\dim(\mathcal{H})/2}^{1}(-k)\}$
$\{v_{1}^{1}(k),\cdots,-Tv_{n}^{1}(-k),v_{n+1}^{1}(k),\cdots,v_{\dim(\mathcal{H})/2}^{1}(k),-Tv_{n+1}^{1}(-k),\cdots-Tv_{\dim(\mathcal{H})/2}^{1}(-k)\}.$
where the added vectors span respectively $\ker(P_{0}(k))$ and $\ker(P_{1}(k))$ and have full periodicity with exceptions
$v_{n+1}^{0}(\pi,k_{2})=e^{i\pi h(\delta)k_{2}}v_{n+1}^{0}(-\pi,k_{2}),\quad v_{n+1}^{1}(\pi,k_{2})=e^{i\pi h(\delta)k_{2}}v_{n+1}^{1}(-\pi,k_{2}).$
Instead, if $\dim(\mathcal{H})=\infty$ we can use the lemma 6.1 to extend the previous frames to
$\{v_{1}^{0}(k),\cdots,-Tv_{n}^{0}(-k),v_{n+1}^{0}(k),-Tv_{n+1}^{0}(-k),\cdots,v_{N}^{0}(k),-Tv_{N}^{0}(-k),\cdots\}$
$\{v_{1}^{1}(k),\cdots,-Tv_{n}^{1}(-k),v_{n+1}^{1}(k),-Tv_{n+1}^{1}(-k),\cdots,v_{N}^{1}(k),-Tv_{N}^{1}(-k),\cdots\}$
where the added vectors have full periodicity and are two orthonormal bases of $\ker(P_{0}(k))$ and $\ker(P_{1}(k))$, respectively.
In both cases, the unitary-valued map $V(k)$ we are looking for is the one such that $V(k)v_{j}^{0}(k)=v_{j}^{1}(k)$ and $V(k)Tv_{j}^{0}(-k)=Tv_{j}^{1}(-k)$ for all $j\in\{1,\cdots,\dim(\mathcal{H})/2\}$ and $k\in\mathbb{T}^{2}$. In fact, it is trivial to prove that $P_{1}(k)=V(k)P_{0}(k)V(k)^{-1}$ and that TV(k)=V(k)TTV(k)=V(-k)T. It is a bit less trivial to prove that this map is periodic despite the fact that it is defined over a non-periodic orthonormal basis. However, we can check, for example, that:
$\begin{split}v_{1}^{1}(\pi,k_{2})&=V(\pi,k_{2})v_{1}^{0}(\pi,k_{2})=V(\pi,k_{2})e^{ih(\delta)k_{2}}v_{1}^{0}(-\pi,k_{2})=e^{ih(\delta)k_{2}}V(\pi,k_{2})v_{1}^{0}(-\pi,k_{2})\\ &=e^{ih(\delta)k_{2}}v_{1}^{1}(-\pi,k_{2})=e^{ih(\delta)k_{2}}V(-\pi,k_{2})v_{1}^{0}(-\pi,k_{2})\end{split}$
so we have $V(-\pi,k_{2})v_{1}^{0}(-\pi,k_{2})=V(\pi,k_{2})v_{1}^{0}(\pi,k_{2})$. Similar arguments can be made for all non-periodic vectors to prove that $V(-\pi,k_{2})$ acts as $V(\pi,k_{2})$.
Now we want to solve the homotopy problem, so if $\delta(P_{0})=\delta(P_{1})$, we want to exhibit a symmetric homotopy between the two projection-valued maps. From the previous point we can create a symmetric unitary equivalence $P_{1}(k)=V(k)P_{0}(k)V(k)^{-1}$ with TV(k)=V(k)TTV(k)=V(-k)T. The first step is to define a unitary-valued map $V_{0}:\mathbb{T}^{2}\to\mathcal{U}(\mathcal{H})$ such that $TV_{0}(k)=V_{0}(-k)T$ and $P_{0}(k)=V_{0}(k)P_{0}(k)V_{0}(k)^{-1}$. If $\dim(\mathcal{H})=\infty$ we can define $V_{0}(k)\equiv\mathds{1}$. Instead, if $\dim(\mathcal{H})=2N<\infty$, we need to do something more. First, we need to use lemma A.1 to select a basis on which $T$ acts as $J\mathcal{K}$. Then we need to observe that the condition TV(k)=V(k)TTV(k)=V(-k)T means that $J\overline{V(k)}=V(-k)J$. Set $\lambda(k):=\det(V(k))$, and consider the triangle $\triangle$ inside $\pi,\pi]^{2}$ with vertices
$Q_{1}=(-\pi,\pi),\quad Q_{2}=(\pi,\pi),\quad Q_{3}=(\pi,-\pi).$
Then on the points $k_{\star}$ fixed by the involution $k\mapsto-k$ we have that $J=V(k_{\star})^{t}JV(k_{\star})$, so $V(k_{\star})\in\mathrm{Sp}(2n)$. Moreover, on the segments $\overline{Q_{1}Q_{2}},\overline{Q_{1},Q_{3}},\overline{Q_{2}Q_{3}}$, we have that $\lambda(k)=\overline{\lambda(-k)}$. Therefore, in those segments, we can apply Lemma A.3 to have that
$\lambda(t,\pi)]_{t\in[-\pi,\pi]}=2m,\quad[\lambda(\pi,t)]_{t\in[-\pi,-\pi]}=2l,\quad[\lambda(t,-t)]_{t\in[-\pi,\pi]}=2h.$
Since the map is defined also inside the triangle, the winding number of $\lambda$ in $\triangle$ must be zero, so $2m=2l+2h$. Now we can use the quasi-periodic frame $\{v_{1}^{0}(k),\cdots,-Tv_{N}^{0}(-k)\}$ to define the unitary operator $V_{0}(k)$ with:
$V_{0}(k)v_{j}^{0}(k)=\begin{cases}e^{i(mk_{1}+lk_{2})}v_{1}^{0}(k)&\mbox{if}\quad j=1\\ v_{j}^{0}(k)&\mbox{otherwise}\end{cases}$
$V_{0}(k)(-Tv_{j}^{0}(-k))=\begin{cases}e^{i(mk_{1}+lk_{2})}(-Tv_{1}^{0}(-k))&\mbox{if}\quad j=1\\ -Tv_{j}^{0}(-k)&\mbox{otherwise}\end{cases}$
This is actually periodic in $k$ because, for example,
$\begin{split}V_{0}(-\pi,k_{2})v_{1}^{0}(\pi,k_{2})&=V_{0}(-\pi,k_{2})e^{i\pi h(\delta)}v_{1}^{0}(-\pi,k_{2})=e^{i\pi h(\delta)}e^{i(-m\pi+lk_{2})}v_{1}^{0}(-\pi,k_{2})=\\ &=e^{-2im\pi}e^{i(m\pi+lk_{2})}e^{i\pi h(\delta)}v_{1}^{0}(-\pi,k_{2})=e^{i(m\pi+lk_{2})}v_{1}^{0}(\pi,k_{2})=\\ &=V_{0}(\pi,k_{2})v_{1}^{0}(\pi,k_{2})\end{split}$
In both cases $\{\dim(\mathcal{H})=\infty,\dim(\mathcal{H})=2N\}$ we want to define a homotopy $V_{t}$ between $V_{0}$ and $V$. To do so, we need to work on $\triangle$. Notice that on the sides of the triangle $\overline{Q_{1}Q_{2}}$, $\overline{Q_{1}Q_{3}}$ and $\overline{Q_{2}Q_{3}}$ the projection-valued map behaves as a one-dimensional time-reversal symmetric projection-valued map:
$TP(k_{1},\pi)=P(-k_{1},-\pi)T=P(-k_{1},\pi)T,$
$TP(\pi,k_{2})=P(-\pi,-k_{2})T=P(\pi,-k_{2})T,$
$TP(k_{1},-k_{1})=P(-k_{1},k_{1})T.$
So we can use Theorem 2.2 to build $V_{t}(k)$ for $k\in\partial\triangle$ with $TV_{t}(k)=V_{t}(-k)T$ for all $t\in[0,1],k\in\partial\triangle$. This means that $V_{t}(k)$ is defined for $(t,k)\in\partial([0,1]\times\triangle)$. This region however, is homeomorphic to $S^{2}$ and since $\pi_{2}(U(2N))=\{0\}=\pi_{2}(\mathcal{U}(\mathcal{H}))$, thanks to Theorems A.7 and A.8, it is always possible to extend the definition inside for $(t,k)\in[0,1]\times\triangle$. As a final adjustment, we just need to define the map in the complementary triangle by imposing $V_{t}(k)=T^{-1}V_{t}(-k)T$ for $k\in[-\pi,\pi]^{2}\setminus\triangle$. Then $V_{t}(k)$ is continuous, periodic and symmetric and therefore:
$P_{t}(k)=V_{t}(k)P_{0}(k)V_{t}(k)^{-1}$
is the symmetric homotopy we were looking for since $P_{0}(k),V_{0}(k)]=0$. ∎
Remark 6.3.
We want to emphasize the role that symmetry plays in this topological study. It is known that projection-valued maps with equal Chern number are homotopic and unitarily equivalent. However, Remark 4.1 states that time-reversal symmetric projection-valued maps always have a trivial Chern number and should always be unitarily equivalent and homotopic. How can this statement agree with the previous theorem? The answer is simple: If two time-reversal symmetric projection-valued maps have different $\delta$ invariants than any homotopy or unitary equivalence connecting the two must break the symmetry constraints present in Definition 1.4. In this way our approach is consistent with the existing theory about vector bundles.
Appendix A Tools from algebraic geometry
In this appendix, we want to prove and highlight some aspects of quaternionic Hilbert spaces and winding numbers that we used in the main text of the paper.
Lemma A.1.
Given a complex Hilbert space $\mathcal{H}$ and an anti-unitary operator $T$ acting on $\mathcal{H}$ with $T^{2}=-\mathds{1}$, we can interpret the pair $(\mathcal{H},T)$ as a quaternionic Hilbert space $\mathcal{H}_{\mathbb{H}}=(\mathcal{H},T)$. Moreover, if $V\subset\mathcal{H}_{\mathbb{H}}$ is a quaternionic Hilbert subspace, namely a complex Hilbert subspace with $TV=V$, then the following statements are true:
  1. (1)
    $V$ admits a quaternionic basis.
  2. (2)
    As a complex Hilbert subspace, $\dim(V)$ is infinite or even.
  3. (3)
    If $\dim(V)=2L$ as a complex Hilbert space, it is always possible to build a complex basis of the form
    $\{v_{1},\cdots,v_{\dim(V)/2},-Tv_{1},\cdots,-Tv_{\dim(V)/2}\}.$
    This means that on this basis $T$ acts as $\mathcal{K}J$, for $\mathcal{K}$ standard complex conjugation and
    (8) $J=\begin{pmatrix}0&\mathbb{I}_{n}\\ -\mathbb{I}_{n}&0\end{pmatrix}$.
This argument is already present in several works, for example in [graf2013bulk]. We prove it below for the sake of completeness.
Proof.
We can define the set composed of the collection of orthonormal vectors that are orthogonal to their partners multiplied with $T$. Namely, we consider the set:
$Z=\left\{\{v_{a}\}_{a\in J}\subset\mathcal{H}|\left\langle v_{a},v_{b}\right\rangle=\delta_{b,a},\left\langle v_{a},Tv_{b}\right\rangle=0\ \forall a,b\in J\right\}$
This set has a natural partial order given by the inclusion, and every increasing sequence
$\{v_{j}\}_{j\in J_{1}}\subset\cdots\subset\{v_{j}\}_{j\in J_{n}}\subset\cdots$
can be viewed up to changing the indexes as a sequence $J_{1}\subset\cdots\subset J_{n}\subset\cdots$. Then we can consider $J=\cup_{i=1}^{\infty}J_{i}$ and look at $\{v_{j}\}_{j\in J}$, which is clearly an element of $Z$. The fact that $T^{2}=-\mathds{1}$ implies that
$\left\langle v,Tv\right\rangle=\overline{\left\langle Tv,T^{2}v\right\rangle}=-\overline{\left\langle Tv,v\right\rangle}=-\left\langle v,Tv\right\rangle$
so $Z$ is not empty because the set consisting of a single normalized vector is an element of it, and we can apply the Zorn lemma to obtain that there is also a maximal element $\{v_{j}\}_{j\in J}$. This maximal element is such that
$\mathcal{H}=\operatorname{Span}_{\mathbb{C}}\left(\{v_{j}\}_{j\in\{1,\cdots,\dim(\mathcal{H})/2\}}\right)\oplus\operatorname{Span}_{\mathbb{C}}\left(\{Tv_{j}\}_{j\in\{1,\cdots,\dim(\mathcal{H})/2\}}\right).$
In fact, by definition, the two spans intersect only at 0 and if, by contradiction, there is a vector $v$ that cannot be obtained by adding elements of the two spans. Then, up to renormalization, the vector
$\tilde{v}=v-\sum_{j\in J}\left\langle v_{j},v\right\rangle v_{j}-\sum_{j\in J}\left\langle Tv_{j},v\right\rangle Tv_{j}$
is non-zero and orthogonal to both spans. Moreover:
$\begin{split}T\tilde{v}&=Tv-\sum_{j\in J}\overline{\left\langle v_{j},v\right\rangle Tv_{j}}+\sum_{j\in J}\overline{\left\langle Tv_{j},v\right\rangle}v_{j}=\\ &=Tv-\sum_{j\in J}\left\langle Tv_{j},Tv\right\rangle Tv_{j}-\sum_{j\in J}\left\langle v_{j},Tv\right\rangle v_{j}\end{split}$
So also $Tv$ is non-trivial and orthogonal to both spans, meaning that we can add $v$ to the collection $\{v_{j}\}_{j\in J}$ and contradict maximality. ∎
Next, we recall the definition and the basic properties of the winding number.
Definition A.2 (Winding number).
Given a map $f:U(1)\to U(1)$, the winding number, often denoted $[f]$, is an integer number exhibited by the natural isomorphism $\pi_{1}(U(1))\simeq\mathbb{Z}$. It can be computed by considering the natural covering $g:\mathbb{R}\to U(1)$ such that $g(t)=e^{i2\pi t}$. Then, the map $f$ can be lifted to a continuous map $\mu:[-\pi,\pi]\to\mathbb{R}$ such that $g(\mu(t))=f(\pi(t))$. Despite the numerous lifting options, the quantity $f]=\mu(\pi)-\mu(-\pi)\in\mathbb{Z}$ is well defined and depends only on the homotopy class of $f$.
Lemma A.3.
If $\lambda:\mathbb{T}^{1}\to U(1)$ is a continuous map with $\lambda(\pi)=\lambda(-\pi)=\lambda(0)=1$ and such that $\lambda(-k)=\overline{\lambda(k)}$, then $\lambda]\in 2\mathbb{Z}$ is even.
Proof.
We can choose a lift $\mu:[-\pi,\pi]\to\mathbb{R}$ such that $e^{i\mu(k)}=\lambda(k)$ with $\mu(0)=0$. Then the properties in the statement translate to the fact that $\mu(k)=-\mu(-k)$ and $\mu(\pi),\mu(-\pi)\in 2\pi\mathbb{Z}$. So, the winding number will be:
$\lambda]=\frac{\mu(\pi)-\mu(-\pi)}{2\pi}=\frac{2\mu(\pi)}{2\pi}\in 2\mathbb{Z}$
Lemma A.4.
If $\lambda:\mathbb{T}^{1}\to\mathbb{T}^{1}$ is a continuous map with $\lambda(k)=\lambda(-k)$, then $\lambda]=0$.
Proof.
We can construct an explicit homotopy $\lambda_{t}$ for $t\in[0,1]$ between $\lambda$ and the constant function $\lambda_{0}\equiv\lambda(\pi)=\lambda(-\pi)$:
$\lambda_{t}(k)=\begin{cases}\lambda(k)&\mbox{for}\quad k\in[-\pi,(t-1)\pi]\cup[(1-t)\pi,\pi]\\ \lambda[(t-1)\pi]=\lambda[(1-t)\pi]&\mbox{for}\quad k\in[(t-1)\pi,(1-t)\pi]\end{cases}$
In fact, the two branches of the function connect for $k=(t-1)\pi,(1-t)\pi$ and also connect for $k=\pi=-\pi$ since $\lambda(\pi)=\lambda(-\pi)$. This means that $\lambda_{t}(k)$ is continuous. ∎
Corollary A.5.
Consider $U(1)=\left\{e^{i\theta}\in\mathbb{C}\right\}$ as a topological group with the standard product of $\mathbb{C}$. If $f,g:U(1)\to U(1)$ are two smooth maps with $f(1)=g(1)=1$ then the winding number of $f\cdot g$ is the sum of the winding number of $f$ and the one of $g$.
Proof.
If we apply the previous theorem we know that $f\cdot g$ is homotopic to fgf*g and it is well known that $\pi_{1}(U(1),1)$ is isomorphic as a group to $\mathbb{Z}$ with the group isomorphism that maps every function to its winding number (see e.g. [may1999concise, Chapter 1, Section 5]). So, if we denote by $[h]$ the winding number of the function $h$, we have that:
$f\cdot g]=[f*g]=[f]+[g].\qed$
Theorem A.6.
If $J=\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}$, $\mathcal{K}$ is the standard complex conjugation and $\mathrm{Sp}(2n)=\{U\in U(2n)|J\mathcal{K}U=UJ\mathcal{K}\}$, then $\mathrm{Sp}(2n)$ is simply connected.
Proof.
A matrix $U$ is in $\mathrm{Sp}(2n)$ if and only if it can be written in exponential form using $U=e^{iH}$, where $H$ is self-adjoint such that $(JH)^{t}=JH$ thanks to Proposition 3.5 and Corollary 11.10 in [hall2013lie]. Since this relation is linear, we can move $U$ toward the identity with the continuous transformation $U_{t}=e^{itH}$ for $t\in[0,1]$. ∎
Theorem A.7.
If $\mathcal{H}$ is a complex Hilbert space with finite dimension, then $\pi_{0}(\mathcal{U}(\mathcal{H}))\simeq\{0\}$, $\pi_{1}(\mathcal{U}(\mathcal{H}))\simeq\mathbb{Z}$ using the homomorphism that computes the winding number of the determinant and $\pi_{2}(\mathcal{U}(\mathcal{H}))$.
Proof.
We can express $\mathcal{U}(\mathcal{H})$ as $\mathrm{U}(n)$, set of $n\times n$ unitary and complex matrices. Then we can express it as a Cartesian product using the bijective map:
$\begin{matrix}\mathrm{U}(n)&\to&S^{1}\times\mathrm{SU}(n)\\ U&\mapsto&\left(\det(U),\begin{pmatrix}\det(U)^{-1}&0\\ 0&\mathbb{I}_{n-1}\end{pmatrix}U\right)\end{matrix}$
This means that there is a short exact sequence of groups
$\{\mathbb{I}\}\to\mathrm{SU}(n)\to\mathrm{U}(n)\to S^{1}\to\{\mathbb{I}\}.$
A sequence like this induces a long exact sequence
$\cdots\to\pi_{j+1}(S^{1})\to\pi_{j}(\mathrm{SU}(n))\to\pi_{j}(\mathrm{U}(n))\to\pi_{j}(S^{1})\to\pi_{j-1}(\mathrm{SU}(n))\to\cdots$
Since $S^{1}$ admits a universal covering using $\mathbb{R}$, we have that $\pi_{j}(S^{1})$ is trivial for $j<1$. Moreover, using the table containing the values of $\pi_{j}(\mathrm{SU}(n))$ in Appendix A, Section 6, Part VII of [ito1993encyclopedic] we have that:
$\pi_{0}(S^{1})\simeq\pi_{2}(S^{1})\simeq\pi_{0}(\mathrm{SU}(n))\simeq\pi_{1}(\mathrm{SU}(n))\simeq\pi_{2}(\mathrm{SU}(n))\simeq\{0\},$
and also that $\pi_{1}(S^{1})\simeq\mathbb{Z}$ using the homomorphism that computes the winding number of the map. Additional details on this topic can be found in [hatcher2005algebraic] on page 344. ∎
Theorem A.8 (Kuiper’s theorem).
If $\mathcal{H}$ is a real, complex, or quaternionic separable Hilbert space with infinite dimension, then $\mathcal{U}(\mathcal{H})$ is weakly contractible.
The proof can be found in [kuiper1965homotopy].
(A. Ferreri) Dipartimento di Matematica, Sapienza Università di Roma
Piazzale Aldo Moro 5, 00185 Rome, Italy
E-mail address: alessandro.ferreri@uniroma1.it
(D. Monaco) Dipartimento di Matematica, Sapienza Università di Roma
Piazzale Aldo Moro 5, 00185 Rome, Italy
E-mail address: domenico.monaco@uniroma1.it
(G. Peluso) Dipartimento di Matematica, Sapienza Università di Roma
Piazzale Aldo Moro 5, 00185 Rome, Italy
E-mail address: gabriele.peluso@uniroma1.it